### v1

# LexGenius: An Expert-Level Benchmark for Large Language Models in Legal General Intelligence

Abstract
Legal general intelligence (GI) refers to artificial intelligence (AI) that encompasses legal understanding, reasoning, and decision-making, simulating the expertise of legal experts across domains. However, existing benchmarks are result-oriented and fail to systematically evaluate the legal intelligence of large language models (LLMs), hindering the development of legal GI. To address this, we propose LexGenius, an expert-level Chinese legal benchmark for evaluating legal GI in LLMs. It follows a Dimension-Task-Ability framework, covering seven dimensions, eleven tasks, and twenty abilities. We use the recent legal cases and exam questions to create multiple-choice questions with a combination of manual and LLM reviews to reduce data leakage risks, ensuring accuracy and reliability through multiple rounds of checks. We evaluate 12 state-of-the-art LLMs using LexGenius and conduct an in-depth analysis. We find significant disparities across legal intelligence abilities for LLMs, with even the best LLMs lagging behind human legal professionals. We believe LexGenius can assess the legal intelligence abilities of LLMs and enhance legal GI development. Our project is available at https://github.com/QwenQKing/LexGenius.
00footnotetext: * Equal contribution.  $\dagger$} Corresponding authors.
Figure 1: Comparison of LLM and human legal experts shows that humans outperform LLMs in all dimensions.
1 Introduction
“Laws are the expression of the will of the people.”
— Montesquieu
Legal general intelligence is the capacity of general AI to perform with expert-level ability across complex legal contexts (e.g., hard tasks, soft intelligence) (Nasir et al., 2024; Kant et al., 2025; Zhou et al., 2025). It involves the precise interpretation of legal provisions, sound inference based on complex factual scenarios Zhang et al. (2025b); Li et al. (2025e); Shen et al. (2025), the resolution of conflicts among rules from multiple interrelated legal domains, and the ability to make normatively binding judgments Yue et al. (2024a); Zhang et al. (2025d); Luo et al. (2025b) in uncertain and ethically sensitive contexts Kim et al. (2025); Huang et al. (2023); Liu et al. (2025c). Legal GI is not just whether AI knows the law, but whether it can participate in the normative structure of legal systems, opening the door to its integration into legal order.
In recent years, LLMs have demonstrated strong performance across general language tasks Mao et al. (2024); Zheng et al. (2025). This progress has catalyzed a surge of interest in adapting LLMs to the legal domain, aiming to tackle challenges, such as legal question-answering  Su et al. (2025); Fei et al. (2024), case analysis Zhang et al. (2025a); Li et al. (2025a), and judgment prediction Liu et al. (2025b); Xie et al. (2025). To assess the legal reasoning capabilities of LLMs, several benchmarks, such as LegalBench Guha et al. (2023), LexEval Li et al. (2024a), and LexGLUE Chalkidis et al. (2021); Jia et al. (2025), are introduced. These benchmarks provide a critical foundation for evaluating, improving, and advancing the legal capabilities of large language models  Luo et al. (2025a, c).
However, the existing benchmarks encounter the following limitations: (1) Legal intelligence has not yet entered the second half of AI. Current benchmarks Chalkidis et al. (2021); Fei et al. (2024) focus on technical tasks while neglecting soft legal intelligence, such as ethical judgment, the law–morality boundary, and societal impact assessment Wang et al. (2024); Cambria et al. (2024). (2) Data contamination and lack of comprehensiveness. Static, publicly available legal benchmarks risk data leakage (Wu et al., 2025) and fail to assess models on dynamic reasoning or novel legal scenarios, leading to overstated and unreliable evaluations. (3) Lack of a structured framework for comprehensively assessing legal intelligence abilities. Outcome-focused benchmarks overlook legal reasoning stages, blurring the line between true understanding and pattern mimicry Cui et al. (2023); Hassani et al. (2025); Chen et al. (2024).
To address the above limitations, we propose LexGenius, a comprehensive benchmark to assess legal general intelligence for LLMs. First, we rethink the evaluation of legal intelligence for LLMs Thakur et al. (2024). Recognizing that existing benchmarks Li et al. (2024b); Wang et al. (2024); Chang et al. (2024); Xu et al. (2025a); Fei et al. (2024) overlook aspects of legal soft intelligence, our framework explicitly incorporates tests of capabilities such as ethical judgment, moral-legal boundaries, and social impact. We have developed a new collection of 8,385 standardized legal multiple-choice questions (MCQs), covering civil, criminal, and commercial law. To ensure legal accuracy, all questions and answers are refined through professional review. These MCQs assess multi-dimensional competencies Cambria et al. (2024); Corfmat et al. (2025) relevant to legal intelligence. Furthermore, to address cognitive coverage limitations in existing benchmarks, we propose a framework structured around the dimensions of legal theory and practice, organizing tasks and abilities to reflect real-world legal intelligence Gursoy and Cai (2025). Focusing on Chinese laws ensures a robust and meaningful assessment, as distinct legal systems would otherwise dilute the evaluation.
Leveraging LexGenius, we evaluated 12 state-of-the-art (SOTA) LLMs and 2 prompting strategies, including naive and chain-of-thought (CoT) prompting Kojima et al. (2022). A baseline performance, constructed by 6 legal professionals, was also established for comparison. Results show that even the top-performing LLM, DeepSeek-R1 Bi et al. (2024), exhibits a significant gap compared to human experts across various legal general intelligence abilities Yao et al. (2025b); Hannah et al. (2025); Dong et al. (2025), as shown in Figure 1. In summary, our contributions in this work include:
  • We propose the LexGenius, a three-level evaluation framework (Dimension-Task-Ability), for systematically and comprehensively evaluating the legal intelligence capabilities of LLMs.
  • We introduce legal soft intelligence into the legal intelligence evaluation of LLMs, paving the way for the assessment of legal general intelligence to move towards the second half of AI.
  • We evaluate 12 SOTA LLMs on LexGenius and analyze their gaps and limitations in legal intelligence at different levels and perspectives.
Benchmark Lan. M-Dim. F-Gra. Com. Soft Int. Ato. Abi.
STARD CN
LexGLUE EN
LegalBench EN
LeCaRDv2 CN
JEC-QA CN
LawBench CN
Legal CQA EN
LexEval CN
Laiw CN
Ours CN
Table 1: Comparison of the existing benchmarks and LexGenius (ours). Lan. means Language; M-Dim. means Multi-Dimensional; F-Gra. means Fine-Grained; Com. means Comprehensiveness; Soft Int. means Soft Intelligence; and Ato. Abi. means Atomicized Ability.
Figure 2: LexGenius can be divided into 3 levels: The first level includes Dimensions 1-7, the second level includes Tasks 1-11, and the third level includes Abilities 1-20 (A. 1 to A. 20). Each is numbered for reference in the text.
2 Related Work
We review existing benchmarks for LLMs, including legal benchmarks and expert-level benchmarks:
Legal Benchmarks. Recently, a series of legal benchmarks have emerged (see Table 1). They have made significant contributions to evaluating LLMs’ performance Kanapala et al. (2019); Yao et al. (2025a), including retrieval (STARD, LeCaRD) Su et al. (2024); Li et al. (2024c), question answering (JEC-QA, Legal CQA) Zhong et al. (2020); Askari et al. (2022), classification (LexGLUE) Chalkidis et al. (2021), reasoning (LegalBench, LexEval) Guha et al. (2023); Li et al. (2024a), and others (Laiw and LawBench) Dai et al. (2025); Fei et al. (2024). However, most benchmarks remain task-oriented and outcome-focused, offering limited insight into the underlying legal general intelligence of LLMs  Yue et al. (2024b).
Expert-level Benchmarks. To usher in the second half of AI, a series of expert-level benchmarks for evaluating LLMs have emerged across various domains Cao et al. (2025); Ni et al. (2025); Li et al. (2025c, d): PhysBench Chow et al. (2025) and PhysReason Zhang et al. (2025c) enhance LLMs’ understanding of physics; MedXpertQA Zuo et al. (2025) and MedagentsbenchTang et al. (2025) focus on medical knowledge; UGMathBench Xu et al. (2025b) assesses math reasoning; and UniToMBench Thiyagarajan et al. (2025) improves theory of mind. Benchmarks like ShotBench Liu et al. (2025a), FinTMMBench Zhu et al. (2025), and Chengyu-Bench Fu et al. (2025) evaluate other fields. However, an expert-level benchmark for legal intelligence is absent.  Wang et al. (2025).
3 LexGenius Framework
In this section, we outline the LexGenius framework (including seven dimensions, eleven tasks, and twenty abilities), which is shown in Figure 2.
3.1 Dimension: Education and Career Drive
The seven legal dimensions of LexGenius are based on Bloom’s Taxonomy of Educational Objectives Bloom et al. (1956), covering the cognitive hierarchy of remembering, understanding, applying, analyzing, evaluating, and creating, alongside the modular model used in legal evaluations across countries, focusing on normative understanding, rule application, procedural operation, and value judgment Wu and Chan (2012); Moon (2020); Parsons et al. (2024). In the hierarchy, remembering and understanding correspond to legal understanding, applying to legal application, analyzing to legal reasoning, evaluating to legal ethics and law and society, creating to advanced arguments, legal language to clarity, and judicial practice to procedural integrity, forming a framework aligned with cognitive principles and professional needs.
Figure 3: The MCQ construction workflow of the LexGenius, which is a process where LLM and manual work are combined. It includes three steps: data collection and structuring, construction of MCQs, and manual review.
3.2 Task: Theory and Practice Coexist
Further, based on Legal Hermeneutics Leyh (2021) and Problem-Solving Cycle theory Stein (1993), we decompose LexGenius’s 7 legal intelligence dimensions into 11 tasks. These tasks align with common legal practice requirements and focus on textual deconstruction, case adaptation, and procedural implementation. Legal Hermeneutics guides the understanding of provisions, critical analysis of texts, and terminology, ensuring accuracy in interpretation. Problem-Solving Cycle theory simulates legal practice, driving reasoning and application analysis for problem-solving, legal and ethical judgment, moral boundaries for value calibration, case reasoning, judicial procedure understanding for validation, and legal impact and social change review, forming a task system for legal problem-solving.
3.3 Ability: Constructivist Learning-based
Furthermore, based on Constructivist Learning Theory Ariati et al. (2025), we extract twenty atomic legal intelligence abilities from the eleven tasks. The theory shifts from outcome assessment to capturing knowledge paths through cognitive traces, simulating real legal scenarios to ensure that the evaluation reflects true professional abilities while aligning cognitive principles with occupational needs. The hierarchy from dimensions to abilities allows LexGenius to perform a detailed, multi-dimensional assessment of LLMs’ legal intelligence, supporting evaluation and optimization.
Figure 4: Data distribution of LexGenius. Left: the MCQ proportions across different laws and the dimensions, tasks, and abilities. Right: the MCQ proportions of abilities. The PRC: the People’s Republic of China.
4 LexGenius Construction
In this section, we introduce the construction principles, construction workflow, data statistics, and evaluation method of the proposed LexGenius.
4.1 Construction Principles
To avoid data leakage and contamination, LexGenius was built from scratch, using recent Chinese legal exam questions and judgment cases. We avoided reusing existing legal datasets to minimize contamination risks and ensure originality. A structured legal competency framework, developed by experienced legal experts, was used to comprehensively cover core legal intelligence abilities Li et al. (2025f). To ensure benchmark effectiveness, a structured review process was established Mohammadi et al. (2025); Li et al. (2025b). Reviewers are master’s candidates in law, systematically trained and thoroughly familiar with key regulations, case frameworks, and legal reasoning methods.
4.2 Construction Workflow
As shown in Figure 3, the process of the construction workflow for legal QA includes three steps:
Step 1: Data collection and structuring. To ensure that the legal basis for the questions is authentic, authoritative, and semantically complete, we systematically collected the latest legal examination question banks and recent judicial cases and used LLM to clean and process these texts in a standardized manner, including encoding format conversion, removal of redundant punctuation, and paragraph reconstruction, to build a structured legal question bank and corpus. Each text is attached with a unique document identifier, source, and usage rights as metadata and saved in a unified JSON structure to facilitate index calls and traceability management when constructing legal MCQs later.
Step 2: Construction of MCQs. There are two methods for constructing multiple-choice questions, both based on LLM. One is to screen and modify legal examination questions. Questions meeting the defined legal abilities are retained; others are modified using LLM. For questions related to legal soft intelligence Schiff et al. (2025), LLM is used to generate them. The first method selects and modifies questions from the legal question bank: MCQs with a single correct answer are retained. In contrast, those with multiple answers are adapted using LLMs and prompt templates to ensure fairness and difficulty. For LLM-generated MCQs, we designed prompt templates with task constraints, ability descriptions, and examples to guide question generation based on legal cases, ensuring unique answers and a clear legal basis.
Step 3: Manual Review. To ensure legal accuracy and competency alignment, we established a team of 9 master candidates in law and created a review process. Each question undergoes double-blind scoring by 2 independent reviewers. The review dimensions include legal accuracy, reasoning rigor, answer uniqueness, competency alignment, and expression standardization, using a five-point scale. When there is a significant discrepancy (e.g., a difference of more than two points), a third reviewer is brought in for arbitration. Questions are retained only if their average score across all dimensions is no less than 4 points and their total score is in the top 500. The final agreement is 99.2%.
4.3 Data Statistics
After multiple rounds of review, the final version of LexGenius consists of 8,385 high-quality legal MCQs. Each question is stored in a structured JSON format, including fields such as question number, competency label, applicable law, question, options, and answers. All data versions are version-controlled, with change logs recorded during updates. To ensure explainability and accountability, we document the construction, review, and modification history of each question, supporting efficient management. Figure 4 shows the number of MCQs for each ability, covering civil disputes, corporate transactions, administrative litigation, criminal litigation, and constitutional rights.
4.4 Evaluation Method
To evaluate the LLM’s legal intelligence, LexGenius adopts a three-level framework Kahan (2015); Huang et al. (2024). The dimension level categorizes tasks into legal cognition areas, providing insight into performance. The task level breaks dimensions into real-world tasks, testing the application, while the ability level evaluates legal abilities, identifying performance differences. At the ability level, scores $A_{i,j,k}$ are the average correctness of MCQs in each ability, where $A_{i,j,k}=\frac{1}{n_{i,j,k}}\sum_{m=1}^{n_{i,j,k}}C_{i,j,k,m}$, with $C_{i,j,k,m}$ the correctness of the $m$-th MCQ in the $k$-th ability of the $j$-th task in the $i$-th dimension, and $n_{i,j,k}$ the number of MCQs for that ability. At the task level, the task score $T_{i,j}$ is the average of ability scores within the task, calculated as $T_{i,j}=\frac{1}{m_{i,j}}\sum_{k=1}^{m_{i,j}}A_{i,j,k}$, where $m_{i,j}$ is the number of abilities in the $j$-th task of the $i$-th dimension. At the dimension level, the dimension score $D_{i}$ is the average of task scores, expressed as $D_{i}=\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}T_{i,j}$, where $n_{i}$ is the number of tasks in the $i$-th dimension.
5 Experiments and Results
In this section, we analyze the experimental results and answer these research questions (RQs): RQ1: Can LLMs’ legal general intelligence rival human legal experts? RQ2: How mature is LLMs’ legal soft intelligence? RQ3: Do LLMs truly understand legal language? RQ4: Can the enhanced methods of LLMs improve their legal intelligence?
Model Legal Und. Legal Rea. Legal App. Legal Ethics Legal Lan. Law & Soc. Judicial Pra. Avg. Nai. CoT Nai. CoT Nai. CoT Nai. CoT Nai. CoT Nai. CoT Nai. CoT Nai. CoT Human 89.14 88.15 89.14 85.19 87.78 87.41 86.67 87.64 Qwen2.5-1.5B 44.05 43.93 41.01 42.28 35.61 35.83 55.70 56.00 51.98 53.32 59.33 59.80 45.50 46.35 47.60 48.22 Qwen2.5-7B 61.41 60.50 57.42 56.57 45.26 45.19 64.00 63.60 66.08 65.92 66.07 66.27 57.55 57.60 59.11 59.09 Qwen3-4B 53.23 52.43 50.37 50.57 39.05 38.07 60.90 61.00 57.63 57.36 62.20 61.13 51.55 51.75 53.13 53.19 Qwen3-8B 58.19 58.77 52.07 52.72 38.73 37.21 62.90 61.60 63.09 61.93 61.20 60.47 55.35 54.40 55.93 55.30 LLaMA-3.2-1B 29.37 28.33 24.03 25.84 26.47 26.23 44.40 43.40 39.37 39.93 50.27 47.93 33.05 32.25 35.57 34.84 LLaMA-3.2-8B 36.99 35.66 31.38 32.85 33.20 33.40 53.50 53.10 46.38 46.62 55.40 56.47 42.80 43.55 42.52 43.09 GLM-4-9B 52.85 53.05 47.78 47.54 35.52 36.47 61.40 61.40 62.46 63.24 61.87 61.47 50.55 51.05 53.49 53.75 DeepSeek-7B 37.27 33.56 34.28 30.96 29.54 30.88 45.10 43.30 39.53 40.79 47.00 47.40 37.20 36.55 38.56 37.63 DeepSeek-R1 67.35 67.76 61.18 61.50 45.91 45.90 70.80 70.60 71.10 71.20 72.40 72.33 62.40 62.85 64.45 64.88 DeepSeek-V3 67.04 67.55 60.90 61.57 46.36 46.60 70.30 69.40 70.48 70.52 71.73 72.60 62.40 62.50 64.17 64.68 GPT‑4o mini 49.77 49.80 43.26 43.91 36.99 36.44 62.20 62.20 57.51 57.53 65.47 65.80 52.25 52.60 52.21 52.32 GPT-4.1 nano 46.51 54.62 43.23 43.88 30.92 40.05 56.80 59.60 53.11 62.69 58.80 60.53 48.20 52.60 48.80 53.00
Table 2: Comparison of Naive (Nai.) and CoT prompts of LLMs on LexGenius (all values in %). Bold entries are the best results with the Naive (CoT) prompt; Underlined entries are the 2nd-best with the Naive (CoT) prompt. (Legal Und. means Legal Understanding; Legal Rea. means Legal Reasoning; Legal App. means Legal Application; Legal Lan. means Legal Language; Law & Soc. means Law and Society; and Judicial Pra. means Judicial Practice.)
5.1 Experimental Setup
We evaluated twelve SOTA LLMs with LexGenius, which include DeepSeek-LLM-7B-Chat (DeepSeek-7B) Bi et al. (2024), Qwen-2.5-7B-Instruct (Qwen-2.5-7B) Hui et al. (2024), Qwen-2.5-1.5B-Instruct (Qwen-2.5-1.5B) Hui et al. (2024), Qwen-3-8B Yang et al. (2025), Qwen-3-4B Yang et al. (2025), GLM-4-9B-Chat (GLM-4-9B) GLM et al. (2024), LLaMA-3.2-1B-Instruct (LLaMA-3.2-1B) Grattafiori et al. (2024), LLaMA-3.2-8B-Instruct (LLaMA-3.2-8B) Grattafiori et al. (2024), DeepSeek-R1 Guo et al. (2025), DeepSeek-V3 Liu et al. (2024), GPT-4o mini Hurst et al. (2024), and GPT‑4.1 nano Brown et al. (2020). We followed the official protocols, using official APIs or LLM weights where applicable. The evaluation utilized two types of prompts: the first type was the Naive prompt; the second was the CoT prompt, encouraging the LLMs to perform step-by-step reasoning. To prevent potential bias, we shuffled the answer options twice and averaged the scores of each LLM.
Model Task 1 Task 2 Task 3 Task 4 Task 5 Task 6 Task 7 Task 8 Task 9 Task 10 Task 11 Avg.
Human 89.14 89.14 89.14 85.19 87.78 87.78 86.67 91.85 82.96 90.00 83.33 87.54
Qwen2.5-1.5B 44.05 41.01 35.61 59.00 52.40 51.98 57.20 56.40 64.40 44.60 46.40 50.28
Qwen2.5-7B 61.41 57.42 45.26 64.80 63.20 66.08 71.00 58.00 69.20 55.20 59.90 61.04
Qwen3-4B 53.23 50.37 39.05 58.00 63.80 57.62 69.00 55.60 62.00 49.50 53.60 55.62
Qwen3-8B 58.19 52.07 38.73 64.20 61.60 63.09 67.60 54.80 61.20 55.50 55.20 57.47
LLaMA-3.2-1B 29.37 24.03 26.47 46.80 42.00 39.37 45.80 46.00 59.00 33.90 32.20 38.63
LLaMA-3.2-8B 36.99 31.38 33.20 56.20 50.80 46.38 55.00 51.60 59.60 43.90 41.70 46.07
GLM-4-9B 52.85 47.78 35.52 66.00 56.80 62.45 64.40 57.60 63.60 49.70 51.40 55.28
DeepSeek-7B 37.27 34.28 29.54 48.40 41.80 39.53 51.60 40.20 49.20 36.90 37.50 40.57
DeepSeek-R1 67.35 61.18 45.91 67.60 74.00 71.09 76.80 65.40 75.00 60.60 64.20 66.29
DeepSeek-V3 67.04 60.90 46.36 67.60 73.00 70.47 76.40 66.20 72.60 60.60 64.20 65.94
GPT‑4o mini 49.77 43.26 36.99 65.20 59.20 57.51 67.60 61.20 67.60 50.30 54.20 55.71
GPT-4.1 nano 46.51 43.23 30.92 59.20 54.40 53.10 63.00 54.40 59.00 48.00 48.40 50.92
Table 3: Performance of twelve LLMs and human experts on eleven legal tasks, showing a significant gap between LLMs and humans. DeepSeek-R1 and DeepSeek-V3 are the top performers, with the greatest challenge in Task 3.
Model A.10 A.11 A.13 A.14 A.15 A.16 Avg.
Human 86.7 83.7 85.9 87.4 91.9 83.0 86.4
Qwen2.5-1.5B 59.0 52.4 60.4 57.2 56.4 64.4 58.3
Qwen2.5-7B 64.8 63.2 68.6 71.0 58.0 69.2 65.8
Qwen3-4B 58.0 63.8 64.6 69.0 55.6 62.0 62.2
Qwen3-8B 64.2 61.6 65.2 67.6 54.8 61.2 62.4
LLaMA-3.2-1B 46.8 42.0 51.0 45.8 46.0 59.0 48.4
LLaMA-3.2-8B 56.2 50.8 63.4 55.0 51.6 59.6 56.1
GLM-4-9B 66.0 56.8 66.2 64.4 57.6 63.6 62.4
DeepSeek-7B 48.4 41.8 46.8 51.6 40.2 49.2 46.3
DeepSeek-R1 67.6 74.0 68.0 76.8 65.4 75.0 71.1
DeepSeek-V3 67.6 73.0 67.4 76.4 66.2 72.6 70.5
GPT-4o mini 65.2 59.2 69.2 67.6 61.2 67.6 65.0
GPT‑4.1 nano 59.2 54.4 60.4 63.0 54.4 59.0 58.4
Table 4: Comparison of the twelve SOTA LLMs for legal soft intelligence on LexGenius. (A. means Ability.)
5.2 Main Results (RQ1)
The performance of the twelve SOTA LLMs across seven dimensions (see Table 2 and Figure 5), eleven tasks (see Table 3), and twenty ability rankings (see Figure 6) on LexGenius is reported.
Comparison with Human. As shown in Table 2 and Figure 5, although LLMs excel in generating legal texts, their capabilities across the seven dimensions still fall short compared to human experts, particularly in areas like legal reasoning, judicial practice, and legal ethics, where value judgments and contextual trade-offs are key. This underscores that legal intelligence is not just about reciting rules but about making sound judgments amidst uncertainty, relying on human experiences, ethical intuition, and institutional understanding. While LLMs can articulate legal principles, they are not yet capable of rendering nuanced judgments. They are powerful assistants, not true counterparts.
Task-view performance. As shown in Table 3, LLMs perform relatively well in static knowledge-based tasks (e.g., legal provisions understanding). However, they are significantly weaker than human experts in tasks that require dynamic reasoning and institutional understanding (e.g., legal application analysis and case reasoning and judgment). Particularly in tasks involving value trade-offs (e.g., legal and ethical judgment), LLMs tend to avoid complex judgments and lack critical thinking and contextual sensitivity. This indicates that they still lack the comprehensive judgment capabilities required for legal practice and remain tools for assistance rather than equivalent intelligent agents.
Figure 5: Comparison of the 12 SOTA LLMs with human experts on 7 core dimensions of legal intelligence.
Ranking of LLMs. Figure 6 reveals that the LLMs’ average scores and rankings are nearly identical. Only a few leading models perform comprehensively and stably, while most rank lower with similar capabilities. This head convergence and tail dispersion pattern suggests that current large models lack balanced legal general intelligence. Their strengths lie in formalized tasks, but they remain weak in complex abilities requiring cross-dimensional integration, value judgment, or institutional understanding. Even top models approaching human-level performance still lack the deep coupling and contextual adaptability required for legal practice, falling short of experts’ capabilities.
Case study. This case (see Figure 8) highlights LLMs’ limitations in legal reasoning: their judgments rely on surface cues, oversimplifying rights conflicts and ignoring core context. While DeepSeek-R1 anchors personality rights, GPT-4o mini misjudges liability, showing a lack of holistic legal understanding. LLMs remain trapped in decontextualized reasoning, unable to balance norms, facts, and values like experts. The gap lies in contextual balancing, a key aspect of legal intelligence.
Naive prompt vs CoT prompt. While CoT enhances surface-level reasoning in several LLMs (see Table 2), it exposes their limitations in high-level legal intelligence (e.g., application, ethics, and judicial practice). The improvement stops at formal logic, failing to capture the nuanced judgments made by human experts that integrate norms, context, and ethics. Humans navigate complexity with stability, while models remain confined to static knowledge reorganization. This highlights that the gap in legal intelligence lies not in reasoning but in making responsible decisions amidst uncertainty, an area current LLMs have yet to bridge.
Figure 6: Average ranking and average score ranking of the 12 SOTA LLMs in the 20 legal intelligence abilities.
Figure 7: Performance of 12 LLMs across 6 legal language indicators, showing gaps compared to the human baseline. Even the best-performing LLMs (e.g., Deepseek-R1) fall short in mastering legal language. Abi. is Ability.
Figure 8: We utilize an MCQ sample case to evaluate DeepSeek-R1 and GPT-4o mini and present the respective thought processes of both LLMs. The English translation of the original Chinese test sample is on the left.
5.3 Legal Soft Intelligence Analysis (RQ2)
As shown in Table 4, results reveal systematic immaturity in LLMs’ legal soft intelligence: LLMs show significant gaps in higher-order abilities like social change, culture, legal coordination, and law-morality boundaries, with fewer issues in analyzing legal enforcement’s social impact. This reflects deficits in experiential social knowledge and ethical reasoning. Scaling fails to overcome performance ceilings, revealing architectural limits in acquiring moral intuition and judgment from static text.
5.4 Legal Language Mastery Analysis (RQ3)
We evaluated the performance of LLMs across nine legal language abilities. The results (see Figure 7) show LLMs excel at reproducing legal text structure and procedural patterns, performing well on formatted tasks. However, their abilities degrade significantly when faced with ambiguity, conflict, or value trade-offs in real legal reasoning. This gap arises from an inherent limitation: models lack understanding of institutional logic, social context, and ethical goals, relying solely on statistical correlations. As a result, LLMs replicate the form of law without truly understanding it and can assist but not replace the essential normative insight and value judgments in legal decision-making.
5.5 With Different Enhanced Methods (RQ4)
As shown in Table 5, the comparison results (more details are in Appendix E) reveal a triple decoupling phenomenon in LLM legal intelligence: (i) Scale-performance decoupling shows that the Qwen2.5 series outperforms the Qwen3 series, with legal intelligence exhibiting a non-linear relationship with model size, indicating that domain-specific pretraining, not scale, drives legal abilities. (ii) Reasoning paradigm decoupling shows that CoT leads to negative transfer in legal tasks due to deterministic constraints, causing probabilistic exploration and semantic drift in closed-solution tasks. (iii) Optimization strategy decoupling shows that supervised fine-tuning benefits weak LLMs but causes catastrophic forgetting in strong ones, while the RAG method reveals the orthogonality of knowledge retrieval and reasoning. Only GRPO (reinforcement learning) achieves stable improvement by aligning reward and evaluative capacity.
Model Baseline CoT SFT RAG GRPO
Qwen2.5-1.5B 48.53 49.44 51.49 34.66 52.47
Qwen2.5-7B 58.45 57.16 55.86 52.22 55.93
Qwen3-4B 29.32 28.39 51.30 37.67 51.00
Qwen3-8B 27.86 27.79 56.85 45.78 55.04
Table 5: Comparison of four LLMs on LexGenius with the enhanced methods, including CoT, Supervised Fine-Tuning (SFT), Retrieval-Augmented Generation (RAG), and Group Relative Policy Optimization (GRPO).
6 Conclusion
In this work, we propose LexGenius, an expert-level and comprehensive benchmark for evaluating LLMs’ legal general intelligence capabilities. Based on the three-level framework (Dimension–Task–Capability), we assess twelve SOTA LLMs from different perspectives. LexGenius addresses gaps in existing benchmarks, including systematic evaluation and alignment with real-world legal reasoning. Experimental results reveal significant legal intelligence gaps of LLMs, highlighting disparities with human legal experts and their specific weaknesses in legal general intelligence.
Appendix
Appendix A Motivation and Theoretical basis of LexGenius
In this section, we primarily explain the design motivations of the LexGenius.
A.1 Design motivation
In this section, we mainly introduce the motivation and reasons for designing LexGenius and answer the question: Why is a structured legal general intelligence evaluation framework needed?
Legal general intelligence is not a stack of tasks, but a simulation of cognitive collaborative chains. Existing evaluation frameworks for legal intelligence often operate at the level of classification or question-answering tasks, presenting only the model’s macro-level accuracy on isolated benchmarks. Such metrics fail to uncover where a model goes wrong within the complex cognitive chain of legal decision-making. Legal intelligence is not a mere sum of discrete tasks—it is an organic coordination of systematic cognitive capabilities, spanning multiple stages such as statutory interpretation, fact extraction, rule adaptation, ethical judgment, and social impact assessment. The performance of an LLM should therefore be decomposed into a multi-stage flow of abilities, rather than treated as a monolithic output. Accordingly, an effective evaluation framework must faithfully map this cognitive chain, avoiding the pitfall of reducing legal intelligence to mere task-solving while overlooking how the model actually thinks.
From performance reporting to capability diagnosis and explanation. Traditional metrics such as accuracy or F1 score can indicate whether a model’s output is right or wrong, but they fail to address a more fundamental question: At which cognitive stage did the model fail? Was it a semantic misunderstanding? A misapplication of legal rules? Or an ethical blind spot? LexGenius introduces 20 atomic legal general intelligence abilities and establishes an interpretable, traceable, and auditable evaluation system through a tri-level mapping mechanism across the ability layer, task layer, and cognitive dimension. This structure not only reveals the model’s deficiencies in specific micro-level legal abilities but also provides an actionable feedback loop for subsequent capability-oriented training, prompt optimization, and safety enhancement.
Enabling cross-stage cognitive analysis and transferability research. The complexity of legal reasoning lies in its chained structure: statutory semantics $\rightarrow$ case fact abstraction $\rightarrow$ rule application $\rightarrow$ ethical judgment $\rightarrow$ precedent alignment. An evaluation framework that fails to distinguish performance across these stages cannot effectively support research in multi-hop reasoning, chain-of-thought attention, or multi-task learning. Through hierarchical abstraction and modular decomposition, our framework standardizes this cognitive pathway, offering a clear experimental baseline for subsequent investigations into how models learn to transfer knowledge and whether they exhibit generalizable reasoning capabilities.
Toward a professional-grade legal general intelligence evaluation paradigm. As LLMs begin to approach the professional thresholds of bar examinations and real-world legal practice, evaluation frameworks must likewise advance to a professional-grade level. LexGenius draws on standards from the National Legal Professional Qualification Examination and international bar exams, distinguishing among multiple specialized dimensions of legal competence, such as legal semantic understanding, norm alignment, ethical judgment, and procedural compliance. This framework breaks away from the general-purpose perspective of traditional NLP benchmarks, aligning instead with the practical demands of legal work. Such a professional, ability-oriented, structured evaluation not only reveals the current boundaries of model capabilities but also illuminates their potential trajectories of future development.
A.2 Framework hierarchy and implementation
The structural design of the LexGenius is illustrated in Table 6. This framework supports both vertical capability dissection—capturing a model’s progressive performance across stages such as legal language understanding → case application → judgment prediction—and horizontal comparison, such as evaluating differences between LLM A and LLM B along the dimension of ethical judgment.
A.3 Toward Legal Cognitive Modeling
Across various domains, an increasing number of expert-level benchmarks are emerging to advance the development and understanding of LLMs. In the legal field, the need for a benchmark that rigorously evaluates expert-level legal general intelligence is equally critical. The proposed three-tier structure—Dimension–Task–Ability—functions not only as an evaluation framework but also as a cognitive modeling paradigm. We move beyond merely assessing outcomes to examining whether a model can think, interpret, and judge like a trained legal professional when confronted with legal contexts. In this sense, LexGenius is not just a benchmark—it is a foundation designed to drive the evolution of legal general intelligence.
Appendix B Definitions of Legal Intelligence Abilities
This section provides detailed definitions for the 20 atomic legal intelligence abilities in the LexGenius framework, following the ability names used in Figure 2. Each ability represents a measurable unit of legal general intelligence, assessed through standardized multiple-choice questions.
1. Precise understanding of legal provisions. Ability to accurately interpret key terms, conditions, and structural logic of legal clauses, including scope and applicability. An MCQ sample of this ability in the LexGenius is shown in Figure 9.
Figure 9: The MCQ sample of ability 1. The left is the original text, and the right is the English translation.
2. Contextual understanding of legal provisions. Ability to interpret legal text within the correct legal and social context, avoiding misinterpretation based on literal reading alone. An MCQ sample of this ability in the LexGenius is shown in Figure 10.
Figure 10: The MCQ sample of ability 2. The left is the original text, and the right is the English translation.
3. Understanding of legal provisions and social phenomena. Ability to relate legal provisions to real-world events, social needs, and historical developments. An MCQ sample of this ability in the LexGenius is shown in Figure 11.
Figure 11: The MCQ sample of ability 3. The left is the original text, and the right is the English translation.
4. Logical ability to reason toward legal conclusions. Ability to construct sound legal arguments based on facts and rules, forming consistent and well-structured conclusions. An MCQ sample of this ability in the LexGenius is in Figure 12.
Figure 12: The MCQ sample of ability 4. The left is the original text, and the right is the English translation.
5. Making reasonable inferences from unclear legal texts. Ability to infer appropriate meanings from vague, ambiguous, or abstract legal language using legal logic and principles. An MCQ sample of this ability in the LexGenius is in Figure 13.
Figure 13: The MCQ sample of ability 5. The left is the original text, and the right is the English translation.
6. Adjusting legal reasoning based on different legal contexts. Ability to adapt reasoning strategies when applying different branches of law, such as civil, criminal, or administrative. An MCQ sample of this ability in the LexGenius is shown in Figure 14.
Figure 14: The MCQ sample of ability 6. The left is the original text, and the right is the English translation.
7. Analyze legal cases. Ability to identify relevant facts and legal issues in a case and link them with the applicable legal norms or precedents. An MCQ sample of this ability in the LexGenius is shown in Figure 15.
Figure 15: The MCQ sample of ability 7. The left is the original text, and the right is the English translation.
8. Choosing and correctly citing the relevant laws. Ability to select the most appropriate legal provisions for a given scenario and cite them accurately in reasoning. An MCQ sample of this ability in the LexGenius is shown in Figure 16.
Figure 16: The MCQ sample of ability 8. The left is the original text, and the right is the English translation.
9. Integrate laws across different fields. Ability to synthesize norms from multiple legal domains and resolve inter-norm conflicts through comprehensive analysis. An MCQ sample of this ability in the LexGenius is shown in Figure 17.
Figure 17: The MCQ sample of ability 9. The left is the original text, and the right is the English translation.
10. Judging the boundary between law and morality and resolving ethical conflicts. Ability to identify and evaluate tensions between legal obligations and moral principles, and propose ethically aware legal judgments. An MCQ sample of this ability in the LexGenius is shown in Figure 18.
Figure 18: The MCQ sample of ability 10. The left is the original text, and the right is the English translation.
11. Critically interpreting legal texts and understanding the lawmakers’ intent. Ability to interpret laws beyond their literal wording by uncovering legislative purpose, background, and systemic coherence. An MCQ sample of this ability in the LexGenius is shown in Figure 19.
Figure 19: The MCQ sample of ability 11. The left is the original text, and the right is the English translation.
12. Interpreting legal terms across fields and adapting to different situations. Ability to understand legal terminology in varied legal contexts and appropriately adapt interpretations to specific domains. An MCQ sample of this ability in the LexGenius is shown in Figure 20.
Figure 20: The MCQ sample of ability 12. The left is the original text, and the right is the English translation.
13. Understanding the exact meaning of legal terms. Ability to grasp the technical definitions, scope, and usage boundaries of domain-specific legal terms. An MCQ sample of this ability in the LexGenius is shown in Figure 21.
Figure 21: The MCQ sample of ability 13. The left is the original text, and the right is the English translation.
14. Analyzing the social impact and stability of legal enforcement. Ability to assess the potential impact of legal implementation on public order, institutional trust, and long-term societal effects. An MCQ sample of this ability in the LexGenius is shown in Figure 22.
Figure 22: The MCQ sample of ability 14. The left is the original text, and the right is the English translation.
15. Social change, culture, and legal coordination. Ability to understand how law responds to social transformation and interacts with culture, economy, and values. An MCQ sample of this ability in the LexGenius is shown in Figure 23.
Figure 23: The MCQ sample of ability 15. The left is the original text, and the right is the English translation.
16. Understanding and managing conflicts between law and morality. Ability to propose socially responsible legal judgments in situations where legal and moral norms collide. An MCQ sample of this ability in the LexGenius is shown in Figure 24.
Figure 24: The MCQ sample of ability 16. The left is the original text, and the right is the English translation.
17. Reasonable legal reasoning and judgment prediction under uncertainty. Ability to make legally sound decisions when faced with ambiguous facts or normative gaps, using analogical reasoning and proportionality. An MCQ sample of this ability in the LexGenius is shown in Figure 25.
Figure 25: The MCQ sample of ability 17. The left is the original text, and the right is the English translation.
18. Case-based reasoning and judgment. Ability to construct judgments through analogical reasoning with relevant precedents and case-specific facts. An MCQ sample of this ability in the LexGenius is shown in Figure 26.
Figure 26: The MCQ sample of ability 18. The left is the original text, and the right is the English translation.
19. Analysis of the application of judicial procedures in different jurisdictions. Ability to identify jurisdictional differences in judicial procedures and adjust legal reasoning accordingly. An MCQ sample of this ability in the LexGenius is shown in Figure 27.
Figure 27: The MCQ sample of ability 19. The left is the original text, and the right is the English translation.
20. Understanding of judicial procedures and the ability to grasp details. Ability to accurately apply procedural rules throughout litigation or non-litigation processes, ensuring procedural compliance. An MCQ sample of this ability in the LexGenius is shown in Figure 28.
Figure 28: The MCQ sample of ability 20. The left is the original text, and the right is the English translation.
Appendix C Experimental setup details
C.1 Large Language Models
We evaluated twelve popular and SOTA LLMs on the proposed LexGenius. For GPT-4o mini, GPT-4 nano, DeepSeek V3, and DeepSeek R1, we accessed the models via their official APIs. For other LLMs, we conducted experimental tests using the official weights. The twelve popular SOTA LLMs are as follows:
Qwen2.5-1.5B-Instruct Hui et al. (2024): A lightweight instruction-tuned model released by Alibaba with 1.5B parameters, designed for edge deployment and local inference with bilingual support and basic task execution.
Qwen2.5-7B-Instruct Hui et al. (2024): A mid-scale model in the Qwen2.5 series, optimized for stronger reasoning and instruction following, suitable for more complex language tasks in medium-sized deployments.
Qwen3-4B Yang et al. (2025): A 4B-parameter model from the third-generation Qwen series, showing strong performance in multilingual, coding, and logical tasks.
Qwen3-8B Yang et al. (2025): An enhanced version of Qwen3 with extended context length and multilingual capabilities, significantly improving performance in complex reasoning and generation tasks.
LLaMA-3.2-1B-Instruct Grattafiori et al. (2024): A compact instruction-tuned model from Meta’s LLaMA 3 series, designed for resource-constrained environments while maintaining core instruction-following capabilities.
LLaMA-3.2-8B-Instruct Grattafiori et al. (2024): A standard model in the LLaMA 3 lineup, offering high-quality multilingual understanding, code generation, and reasoning, achieving state-of-the-art results across many tasks.
GLM-4-9B-Chat GLM et al. (2024): A 9B bilingual chat model developed by Zhipu AI, equipped for multi-turn dialogue, tool use, and contextual memory, with strong performance particularly in Chinese semantic understanding.
DeepSeek-LLM-7B-Chat Bi et al. (2024): A 7B bilingual chat model by DeepSeek, integrating capabilities in code generation, mathematics, and language understanding, suitable for dialogue and multitask settings.
DeepSeek-R1 Guo et al. (2025): A large-scale open-source language model developed by the Chinese company DeepSeek, featuring strong capabilities in mathematics, programming, and reasoning with efficient training and leading performance.
DeepSeek-V3 Liu et al. (2024): A large language model based on a mixture-of-experts architecture, excelling in mathematics, programming, and logical reasoning, and is well-suited for a variety of intelligent application scenarios.
GPT-4o mini Hurst et al. (2024): A parameter-efficient version of OpenAI’s GPT-4o model, supporting multimodal inputs (text, image, audio) with consistent alignment behavior as its full-size counterpart.
GPT-4.1 nano Hurst et al. (2024): An ultra-compact model in the GPT-4.1 series, designed for on-device and embedded inference with support for moderate-length contexts and basic reasoning under resource constraints.
Figure 29: The two utilized prompt methods for LLMs. In this figure, we provide the Chinese and English texts.
C.2 Two prompt methods
We employed two prompting strategies to evaluate the legal general intelligence on the LexGenius. One is a simple prompt (naive), and the other is a Chain-of-Thought (CoT) prompt. These two strategies are illustrated in Figure 29.
Naive Prompt is the most basic form of prompt design, where the task objective is directly stated to the language model, with the expectation that it will generate the final output without revealing any intermediate reasoning steps. This approach relies on the model’s end-to-end learning ability, requiring it to complete the full input-to-output mapping within a single forward pass. Its main advantage lies in simplicity and speed, but it places high demands on the model’s implicit knowledge and generalization capability. Especially for complex tasks, it is prone to errors due to semantic ambiguity or interference from multi-step reasoning chains. Conceptually, this prompting strategy does not explicitly activate the model’s thinking path, which may lead it to overlook latent contextual connections, implicit assumptions, or task structure. As task complexity increases, performance tends to degrade rapidly due to the lack of a mechanism for step-by-step reasoning construction.
CoT Prompt is to explicitly guide the model through step-by-step reasoning by decomposing a task into a sequence of intermediate steps expressed in natural language Kojima et al. (2022). This approach encourages the model to simulate the human problem-solving process and functions as a form of prompt programming, activating internal causal chains and logical pathways by embedding intermediate reasoning states within the prompt. Mechanistically, CoT prompting leverages the structured knowledge and implicit logical modeling capabilities that large models naturally acquire during training. When the prompt includes cues such as think, step, or other structured instructions, the model tends to produce coherent reasoning processes rather than skipping directly to the final answer. This reasoning chain helps reduce error rates in complex tasks and improves the model’s ability to handle multi-step dependencies, conditional logic, and task decomposition.
Appendix D Results of 20 Legal Intelligence Abilities
Based on the structured capability framework provided by LexGenius, we evaluated the performance of various SOTA LLMs across 20 legal intelligence abilities (see Table 7). LexGenius is categorized under 7 core dimensions. These dimensions are further divided into 11 tasks and 20 abilities, covering a comprehensive range of legal intelligence abilities, including understanding, reasoning, application, ethical judgment, language processing, socio-legal interaction, and judicial practice.
The results of 20 legal intelligence abilities for the 12 SOTA LLMs are as illustrated in Table 7. The LLMs’ legal intelligence abilities decline significantly in tasks requiring deeper abstraction. These tasks involve complex value judgments, cross-domain norm integration, and procedural reasoning—areas where LLMs struggle to match human-like legal cognition. This highlights the need for further model optimization in sociological, ethical, and institutional aspects of legal general intelligence.
Level Description Goals and Functions Number Dimension Core legal intelligence focus areas and judgments (e.g., legal reasoning) Build top-level cognition and capability aggregation 7 Task Scenario-based legal tasks (e.g., case reasoning and judgment prediction) Mid-level structure linking abilities and test design 11 Ability Measurable legal intelligence ability units (e.g., statute understanding and interpretation) Minimum unit, supporting fine-grained evaluation and diagnosis 20
Table 6: Hierarchical levels of the LexGenius and corresponding implementation counts. It includes the Dimension level (high-level cognitive targets), the Task level (scenario-based applications), and the Ability level (fine-grained evaluable units), along with the number of implemented benchmarks under each category.
Figure 30: The correlation analysis of legal intelligence ability, task, and dimension in LexGenius for 12 LLMs.
Appendix E With Different Enhanced Methods
To accurately compare and evaluate the impact of different optimization and enhancement methods on the legal intelligence capabilities of Large Language Models (LLMs), we selected four LLMs (including Qwen2.5-1.5B-Instruct, Qwen2.5-7B-Instruct, Qwen3-4B, and Qwen3-8B) and experimented with Supervised Fine-Tuning (SFT), Chain-of-Thought (CoT), Retrieval-Augmented Generation (RAG), and Reinforcement Learning (RL) algorithms. We randomly sampled 64 test instances from each of the 20 ability test sets in LexGenius, resulting in 1,280 total samples for evaluation. The remaining 7,105 data samples were used as the training set for SFT and RL, as well as for constructing the retrieval corpus. The experimental results of these LLMs, after applying these enhancement methods across various dimensions and tasks of legal intelligence, are illustrated in Table 8 and Table 9.
Appendix F Correlation Analysis
The average performance of the twelve SOTA LLMs on LexGenius is utilized to conduct correlation analysis on the benchmark (as shown in Figure 30). It illustrates that most of the legal intelligence abilities (left), tasks (upper right), and dimensions (lower right) exhibit low correlations. It shows the effectiveness of the developed questions in our LexGenius, because the low intercorrelation suggests that models cannot rely on general legal heuristics or shallow transfer across domains to perform well; instead, success in one category does not guarantee success in others. This reflects the comprehensive coverage and conceptual independence of our benchmark dimensions, further validating their robustness as an evaluation framework.
Appendix G Limitation and Future Work
To bridge the gap between current benchmark capabilities and the demands of real-world legal practice, we outline several directions for future work. While LexGenius establishes a foundational framework for evaluating legal general intelligence, it remains constrained by unimodal design, limited jurisdictional scope, and a lack of temporal modeling. To address these limitations, future iterations will pursue enhancements in three key areas: incorporating multimodal legal tasks, expanding linguistic and legal diversity, and introducing dynamic temporal reasoning. These improvements aim to more accurately reflect the complexity of legal environments and to enable more robust, generalizable evaluations of large language models. We hope that LexGenius and these discussions can promote the development and research of legal general intelligence.
G.1 Limitation
While LexGenius offers a structured framework for evaluating legal general intelligence in large language models, it still falls short in key areas that limit its realism and generalizability. Specifically, it lacks support for multimodal inputs, is confined to a single language and legal system, and does not account for the temporal dynamics of law. These gaps constrain its alignment with real-world legal complexity and hinder its ability to fully assess models’ applicability in diverse, evolving legal contexts. The following subsections detail these core limitations.
Lack of Multimodal Tasks Limits Realistic Evidence Modeling. The current version of LexGenius is built and evaluated entirely on pure textual materials, without incorporating multimodal evidence types commonly found in real-world cases, such as scanned contract images, video stills, or audio transcriptions. This unimodal design fails to assess a model’s capabilities in visual perception, auditory understanding, and cross-modal reasoning, which are essential for handling actual judicial cases. The absence of multimodal inputs limits the applicability of LLMs in tasks such as evidence review, fact reconstruction, and the interpretation of visual-legal content, thereby reducing the fidelity of the evaluation to real-world legal scenarios.
Linguistic and Jurisdictional Limitations Undermine Cross-Cultural Generalization. LexGenius is currently constructed solely from Chinese-language corpora and based on the civil law system of Mainland China, exhibiting clear singularity in both linguistic and legal dimensions. As a result, model evaluations are only valid within the context of "Chinese civil law" and fail to capture broader capabilities such as interpreting international statutes, translating case law, or conducting comparative legal analysis. In tasks involving global legal services, cross-border disputes, or foreign compliance, this limitation significantly hampers a model’s transferability and generalization, impeding its evolution into a truly universal system of legal intelligence.
Ability Human LLM 1 LLM 2 LLM 3 LLM 4 LLM 5 LLM 6 LLM 7 LLM 8 LLM 9 LLM 10 LLM 11 LLM 12 Naive Prompt Ability 1 89.63 45.00 58.20 50.80 55.20 28.80 37.80 53.00 39.00 64.40 64.80 47.00 41.40 Ability 2 87.41 37.78 60.89 46.22 54.22 30.67 33.33 46.22 34.22 67.11 65.78 47.11 40.44 Ability 3 90.37 49.38 65.15 62.66 65.15 28.63 39.83 59.34 38.59 70.54 70.54 55.19 57.68 Ability 4 92.59 39.40 56.60 48.00 50.60 25.80 32.20 47.60 32.20 59.20 58.00 46.00 39.00 Ability 5 88.89 43.45 55.36 52.38 50.60 23.81 27.98 41.67 35.71 60.71 60.12 39.29 47.62 Ability 6 85.93 40.19 60.29 50.72 55.02 22.49 33.97 54.07 34.93 63.64 64.59 44.50 43.06 Ability 7 85.93 36.20 47.80 41.40 42.80 25.60 31.60 39.40 35.80 46.80 46.80 40.40 35.40 Ability 8 92.59 39.60 52.20 46.00 48.40 23.20 34.80 42.60 27.40 53.00 55.20 40.40 35.80 Ability 9 88.89 31.03 35.78 29.74 25.00 30.60 33.19 24.57 25.43 37.93 37.07 30.17 21.55 Ability 10 86.67 59.00 64.80 58.00 64.20 46.80 56.20 66.00 48.40 67.60 67.60 65.20 59.20 Ability 11 83.70 52.40 63.20 63.80 61.60 42.00 50.80 56.80 41.80 74.00 73.00 59.20 54.40 Ability 12 89.63 43.55 63.55 50.65 60.97 27.74 29.35 58.71 32.26 74.19 73.55 45.81 45.81 Ability 13 85.93 60.40 68.60 64.60 65.20 51.00 63.40 66.20 46.80 68.00 67.40 69.20 60.40 Ability 14 87.41 57.20 71.00 69.00 67.60 45.80 55.00 64.40 51.60 76.80 76.40 67.60 63.00 Ability 15 91.85 56.40 58.00 55.60 54.80 46.00 51.60 57.60 40.20 65.40 66.20 61.20 54.40 Ability 16 82.96 64.40 69.20 62.00 61.20 59.00 59.60 63.60 49.20 75.00 72.60 67.60 59.00 Ability 17 89.63 38.40 45.80 40.80 45.00 23.80 34.20 39.40 33.40 48.80 49.40 40.20 37.00 Ability 18 90.37 50.80 64.60 58.20 66.00 44.00 53.60 60.00 40.40 72.40 71.80 60.40 59.00 Ability 19 79.26 58.20 74.00 65.80 69.60 37.80 54.60 63.80 45.60 80.00 81.20 65.80 61.80 Ability 20 87.41 34.60 45.80 41.40 40.80 26.60 28.80 39.00 29.40 48.40 47.20 42.60 35.00 CoT Prompt Ability 1 89.63 45.00 57.60 50.60 54.40 30.00 36.40 53.20 36.80 64.40 64.20 48.00 56.80 Ability 2 87.41 38.67 60.00 44.44 55.11 28.44 32.00 45.78 30.67 66.67 66.67 46.22 50.22 Ability 3 90.37 48.13 63.90 62.24 66.80 26.56 38.59 60.17 33.20 72.20 71.78 55.19 56.85 Ability 4 92.59 42.00 56.20 48.00 49.20 23.60 32.20 47.00 32.00 60.40 58.80 45.80 40.00 Ability 5 88.89 42.26 54.17 52.98 50.60 28.57 30.95 41.07 27.38 59.52 61.31 40.48 47.62 Ability 6 85.93 42.58 59.33 50.72 58.37 25.36 35.41 54.55 33.49 64.59 64.59 45.45 44.02 Ability 7 85.93 36.00 48.00 42.00 41.40 24.60 32.00 39.60 34.40 46.60 47.60 39.40 39.40 Ability 8 92.59 39.60 51.80 44.20 47.40 22.20 35.00 41.80 29.80 53.60 53.40 40.60 48.00 Ability 9 88.89 31.90 35.78 28.02 22.84 31.90 33.19 28.02 28.45 37.50 38.79 29.31 32.76 Ability 10 86.67 59.40 65.40 58.60 62.80 46.60 55.20 65.60 50.20 68.00 67.40 65.80 60.80 Ability 11 83.70 52.60 61.80 63.40 60.40 40.20 51.00 57.20 36.40 73.20 71.40 58.60 58.40 Ability 12 89.63 44.84 63.23 50.32 60.65 26.45 29.03 59.68 32.58 74.19 73.23 46.45 60.97 Ability 13 85.93 61.80 68.60 64.40 63.20 53.40 64.20 66.80 49.00 68.20 67.80 68.60 64.40 Ability 14 87.41 57.20 70.40 68.20 66.00 41.40 56.40 63.40 50.40 77.00 77.20 67.20 66.20 Ability 15 91.85 56.20 59.60 54.00 54.60 45.60 51.80 57.40 42.60 65.00 65.20 61.40 54.00 Ability 16 82.96 66.00 68.80 61.20 60.80 56.80 61.20 63.60 49.20 75.00 75.40 68.80 61.40 Ability 17 89.63 38.80 45.20 41.20 44.40 21.60 32.40 38.40 30.60 51.20 48.80 39.60 37.00 Ability 18 90.37 50.80 64.60 58.00 65.00 42.80 54.80 60.80 42.60 73.60 72.80 61.80 60.60 Ability 19 79.26 59.20 74.40 66.00 68.40 39.00 56.60 64.60 43.60 79.60 81.00 67.00 66.00 Ability 20 87.41 36.60 46.20 41.80 39.80 25.60 30.40 40.40 29.40 47.00 47.40 42.00 46.80
Table 7: Comparison of performance across 20 legal intelligence abilities for Naive Prompt and CoT Prompt on various LLMs (all values in %). LLM 1 is Qwen2.5-1.5B-Instruct; LLM 2 is Qwen2.5-7B-Instruct; LLM 3 is Qwen3-4B; LLM 4 is Qwen3-8B; LLM 5 is Llama-3.2-1B-Instruct; LLM 6 is Llama-3.2-8B-Instruct; LLM 7 is GLM-4-9B-Chat; LLM 8 is DeepSeek-LLM-7B-Chat; LLM 9 is DeepSeek-R1; LLM 10 is DeepSeek-V3; LLM 11 is GPT-4o mini; and LLM 12 is GPT-4.1 nano.
Lack of Evaluation on Temporal Sensitivity and Legal Validity Awareness. A core characteristic of law is its temporal nature—the applicable rules for a given issue may vary significantly across time, especially before and after legislative amendments. LexGenius currently does not incorporate a systematic temporal dimension to assess whether models can understand the time-bound applicability of statutes, the validity period of precedents, or transitional legal provisions. Without such temporal sensitivity tests, models may produce outdated or legally invalid answers when facing evolving legal frameworks, with no mechanism to detect these errors.
G.2 Future work
LexGenius aims to advance Chinese legal intelligence toward generalization. The current version covers fundamental legal domains and establishes a Dimension–Task–Ability evaluation framework, yet it still falls short of capturing the full complexity of real-world legal systems. Therefore, our future work will primarily focus on the following:
Incorporating Multimodal Tasks to Enhance Realistic Evidence Modeling. The current version is constructed solely on textual materials and does not include multimodal information commonly encountered in real legal cases, such as scanned contracts, courtroom audio transcriptions, or surveillance video stills. The absence of such inputs limits the evaluation of model capabilities in visual perception, auditory comprehension, and cross-modal reasoning, all of which are essential for key legal tasks such as evidence review, fact reconstruction, and interpretation of visual-legal content. In future iterations, we plan to embed images, audio, and other modalities into tasks to assess models’ ability to reason and judge based on heterogeneous, multi-source information, thus aligning evaluation more closely with practical judicial needs.
Expanding Linguistic and Jurisdictional Coverage to Improve Cross-Cultural Generalization. The current dataset is grounded in Chinese-language texts and the civil law system of Mainland China, exhibiting limitations in both language and legal tradition. This restricts the applicability of evaluation outcomes. Future versions will incorporate legal texts from Hong Kong, Macau, and Taiwan, as well as English-language statutes and case law from common law systems. We aim to construct bilingual legal QA pairs, statute translation tasks, and comparative law analyses to support the evaluation of models’ capabilities in understanding, aligning, and adapting across legal and linguistic contexts. This expansion will contribute to the training and benchmarking of legal language models equipped for global legal services.
Introducing Dynamic Testing of Legal Temporality and Time Sensitivity. Legal applicability is highly time-dependent—amendments to laws can lead to drastically different rulings, and precedents often carry specific periods of validity and applicability. Currently, LexGenius lacks a systematic temporal dimension, making it difficult to evaluate whether a model can identify the applicable time windows of statutes, conditions for transitional provisions, or conflicts between old and new laws. Future versions will include temporally structured legal tasks that require models to make dynamic judgments under varying legal timeframes, thereby enhancing their understanding and adaptability to evolving legal systems.
Model Legal Und. Legal Rea. Legal App. Legal Ethics Legal Lan. Law & Soc. Judicial Pra. Avg.
Baseline
Qwen3-4B 22.39 31.77 32.82 30.47 18.75 33.85 35.16 29.32
Qwen2.5-7B 62.50 51.04 42.19 61.72 63.28 69.79 58.60 58.45
Qwen3-8B 25.52 31.25 29.69 27.34 17.97 32.81 30.47 27.86
Qwen2.5-1.5B 43.75 44.79 35.41 56.25 51.57 58.33 49.61 48.53
CoT
Qwen3-4B 19.79 29.69 34.38 29.69 18.75 34.37 32.03 28.39
Qwen2.5-7B 60.94 52.09 38.54 60.16 61.72 69.27 57.42 57.16
Qwen3-8B 23.43 29.69 33.33 28.91 19.53 30.73 28.91 27.79
Qwen2.5-1.5B 42.19 46.35 35.42 57.82 52.35 61.98 50.00 49.44
RAG
Qwen3-4B 37.67 35.64 34.43 36.86 36.35 42.70 40.02 37.67
Qwen2.5-7B 57.29 45.84 41.14 53.13 54.89 57.77 55.47 52.22
Qwen3-8B 48.96 47.92 28.64 51.56 48.44 50.00 44.93 45.78
Qwen2.5-1.5B 30.93 34.54 32.08 37.50 32.81 38.71 36.06 34.66
SFT
Qwen3-4B 50.52 37.50 33.86 67.19 52.35 67.71 50.00 51.30
Qwen2.5-7B 56.25 52.08 31.25 64.85 61.72 66.66 58.20 55.86
Qwen3-8B 61.98 45.83 32.81 69.53 63.28 71.35 53.13 56.84
Qwen2.5-1.5B 49.48 43.75 31.25 60.16 61.72 67.19 46.87 51.49
GRPO
Qwen3-4B 52.08 47.92 29.17 60.94 53.91 63.02 50.00 51.01
Qwen2.5-7B 57.29 53.65 35.42 56.25 67.97 63.54 57.42 55.93
Qwen3-8B 59.90 50.00 35.94 60.94 63.28 62.50 52.74 55.04
Qwen2.5-1.5B 53.65 46.36 32.81 62.50 60.94 61.46 49.61 52.48
Table 8: Comparison of the four LLMs with different enhanced methods on seven dimensions of LexGenius, which include CoT, RAG, SFT, and GRPO.
Model Task 1 Task 2 Task 3 Task 4 Task 5 Task 6 Task 7 Task 8 Task 9 Task 10 Task 11 Avg.
Baseline
Qwen3-4B 22.39 31.77 32.82 28.12 32.81 18.75 37.50 39.06 25.00 30.46 39.84 30.78
Qwen2.5-7B 62.50 51.04 42.19 59.38 64.06 63.28 68.75 67.19 73.44 53.91 63.28 60.82
Qwen3-8B 25.52 31.25 29.69 28.12 26.56 17.97 35.94 37.50 25.00 27.34 33.59 28.95
Qwen2.5-1.5B 43.75 44.79 35.41 57.81 54.69 51.56 53.12 60.94 60.94 50.00 49.22 51.11
CoT
Qwen3-4B 19.79 29.69 34.38 28.12 31.25 18.75 40.62 37.50 25.00 28.91 35.16 29.92
Qwen2.5-7B 60.94 52.09 38.54 56.25 64.06 61.72 67.19 67.19 73.44 52.34 62.50 59.66
Qwen3-8B 23.43 29.69 33.33 29.69 28.12 19.53 29.69 39.06 23.44 25.00 32.81 28.53
Qwen2.5-1.5B 42.19 46.35 35.42 59.38 56.25 52.34 53.12 65.62 67.19 51.56 48.44 52.53
RAG
Qwen3-4B 37.67 35.64 34.43 38.71 35.00 36.35 54.84 30.65 42.62 40.05 39.98 38.72
Qwen2.5-1.5B 30.93 34.54 32.08 40.62 34.38 32.81 33.33 50.00 32.81 38.28 33.84 35.78
Qwen3-8B 48.96 47.92 28.64 53.12 50.00 48.44 53.12 43.75 53.12 42.19 47.66 46.99
Qwen2.5-7B 57.29 45.84 41.14 56.25 50.00 54.89 55.56 58.06 59.68 50.78 60.16 53.60
SFT
Qwen3-4B 50.52 37.50 33.86 73.44 60.94 52.34 60.94 70.31 71.88 46.09 53.91 55.61
Qwen2.5-7B 56.25 52.08 31.25 70.31 59.38 61.72 65.62 65.62 68.75 59.38 57.03 58.85
Qwen3-8B 61.98 45.83 32.81 75.00 64.06 63.28 64.06 68.75 81.25 53.91 52.34 60.30
Qwen2.5-1.5B 49.48 43.75 31.25 70.31 50.00 61.72 64.06 67.19 70.31 48.44 45.31 54.71
GRPO
Qwen3-4B 52.08 47.92 29.17 62.50 59.38 53.91 67.19 56.25 65.62 49.22 50.78 54.00
Qwen2.5-7B 57.29 53.65 35.42 57.81 54.69 67.97 62.50 60.94 67.19 61.72 53.12 57.48
Qwen3-8B 59.90 50.00 35.94 60.94 60.94 63.28 57.81 59.38 70.31 55.47 50.00 56.72
Qwen2.5-1.5B 53.65 46.36 32.81 68.75 56.25 60.94 62.50 56.25 65.62 52.34 46.87 54.76
Table 9: Comparison of the four LLMs with different enhanced methods on eleven tasks of LexGenius, which include CoT, RAG, SFT, and GRPO.

### v2

@@
 00footnotetext: * Equal contribution.  $\dagger$} Corresponding authors.
-Figure 1: Comparison of LLM and human legal experts shows that humans outperform LLMs in all dimensions.
+Figure 1: Comparison of the state-of-the-art LLMs and human legal experts illustrates that humans outperform LLMs in the seven legal intelligence dimensions.
 1 Introduction
-“Laws are the expression of the will of the people.”
-— Montesquieu
-Legal general intelligence is the capacity of general AI to perform with expert-level ability across complex legal contexts (e.g., hard tasks, soft intelligence) (Nasir et al., 2024; Kant et al., 2025; Zhou et al., 2025).
-It involves the precise interpretation of legal provisions, sound inference based on complex factual scenarios Zhang et al.
-(2025b); Li et al.
-(2025e); Shen et al.
-(2025), the resolution of conflicts among rules from multiple interrelated legal domains, and the ability to make normatively binding judgments Yue et al.
-(2024a); Zhang et al.
-(2025d); Luo et al.
-(2025b) in uncertain and ethically sensitive contexts Kim et al.
-(2025); Huang et al.
-(2023); Liu et al.
-(2025c).
-Legal GI is not just whether AI knows the law, but whether it can participate in the normative structure of legal systems, opening the door to its integration into legal order.
-In recent years, LLMs have demonstrated strong performance across general language tasks Mao et al.
-(2024); Zheng et al.
-(2025).
-This progress has catalyzed a surge of interest in adapting LLMs to the legal domain, aiming to tackle challenges, such as legal question-answering  Su et al.
-(2025); Fei et al.
-(2024), case analysis Zhang et al.
-(2025a); Li et al.
-(2025a), and judgment prediction Liu et al.
-(2025b); Xie et al.
-(2025).
-To assess the legal reasoning capabilities of LLMs, several benchmarks, such as LegalBench Guha et al.
-(2023), LexEval Li et al.
-(2024a), and LexGLUE Chalkidis et al.
-(2021); Jia et al.
-(2025), are introduced.
-These benchmarks provide a critical foundation for evaluating, improving, and advancing the legal capabilities of large language models  Luo et al.
-(2025a, c).
+“The law is the expression of the general will.”
+— Jean-Jacques Rousseau
+Legal general intelligence is the capacity of general AI to perform with expert-level ability across complex legal contexts (e.g., hard tasks, soft intelligence) (kant2025towards; zhou2025lawgpt).
+It involves the precise interpretation of legal provisions, sound inference based on complex factual scenarios zhang2025syler; li2025basis; SHEN2025102860, the resolution of conflicts among rules from multiple interrelated legal domains, and the ability to make normatively binding judgments yue2024circumstance; zhang2025rljp; luo2025hypergraphrag in uncertain and ethically sensitive contexts kim2025legisflow; huang2023lawyer; liu2025prompt.
+Legal general intelligence is not just whether AI knows the law, but whether it can participate in the normative structure of legal systems, thereby opening the door to its integration into legal order.
+In recent years, LLMs have demonstrated strong performance across general language tasks mao2024gpteval; zheng2025towards.
+This progress has catalyzed a surge of interest in adapting LLMs to the legal domain, aiming to tackle challenges, such as legal question-answering  su2025judge; fei2024lawbench, case analysis zhang2025citalaw; li2025unilr, and judgment prediction liu2025legal; xie2025lawchain.
+To assess the legal reasoning capabilities of LLMs, several benchmarks, such as LegalBench guha2023legalbench, LexEval li2024lexeval, and LexGLUE chalkidis2021lexglue; jia2025ready, are introduced.
+These benchmarks provide a critical foundation for evaluating, improving, and advancing the legal capabilities of large language models  luo2025graph; luo2025kbqao1.
 However, the existing benchmarks encounter the following limitations: (1) Legal intelligence has not yet entered the second half of AI.
-Current benchmarks Chalkidis et al.
-(2021); Fei et al.
-(2024) focus on technical tasks while neglecting soft legal intelligence, such as ethical judgment, the law–morality boundary, and societal impact assessment Wang et al.
-(2024); Cambria et al.
-(2024).
+Current benchmarks chalkidis2021lexglue; fei2024lawbench focus on technical tasks while neglecting soft legal intelligence, such as ethical judgment, the law–morality boundary, and societal impact assessment wang2024legal; cambria2024xai.
 (2) Data contamination and lack of comprehensiveness.
-Static, publicly available legal benchmarks risk data leakage (Wu et al., 2025) and fail to assess models on dynamic reasoning or novel legal scenarios, leading to overstated and unreliable evaluations.
+Static, publicly available legal benchmarks risk data leakage (wu2025antileak) and fail to assess models on dynamic reasoning or novel legal scenarios, leading to overstated and unreliable evaluations.
 (3) Lack of a structured framework for comprehensively assessing legal intelligence abilities.
-Outcome-focused benchmarks overlook legal reasoning stages, blurring the line between true understanding and pattern mimicry Cui et al.
-(2023); Hassani et al.
-(2025); Chen et al.
-(2024).
+Outcome-focused benchmarks overlook legal reasoning stages, blurring the line between true understanding and pattern mimicry cui2023chatlaw; hassani2025empirical; chen2024survey.
 To address the above limitations, we propose LexGenius, a comprehensive benchmark to assess legal general intelligence for LLMs.
-First, we rethink the evaluation of legal intelligence for LLMs Thakur et al.
-(2024).
-Recognizing that existing benchmarks Li et al.
-(2024b); Wang et al.
-(2024); Chang et al.
-(2024); Xu et al.
-(2025a); Fei et al.
-(2024) overlook aspects of legal soft intelligence, our framework explicitly incorporates tests of capabilities such as ethical judgment, moral-legal boundaries, and social impact.
+First, we rethink the evaluation of legal intelligence for LLMs thakur2024judging.
+Recognizing that existing benchmarks li2024llms; wang2024legal; chang2024survey; xu2025towards; fei2024lawbench overlook aspects of legal soft intelligence, our framework explicitly incorporates tests of capabilities such as ethical judgment, moral-legal boundaries, and social impact.
 We have developed a new collection of 8,385 standardized legal multiple-choice questions (MCQs), covering civil, criminal, and commercial law.
 To ensure legal accuracy, all questions and answers are refined through professional review.
-These MCQs assess multi-dimensional competencies Cambria et al.
-(2024); Corfmat et al.
-(2025) relevant to legal intelligence.
-Furthermore, to address cognitive coverage limitations in existing benchmarks, we propose a framework structured around the dimensions of legal theory and practice, organizing tasks and abilities to reflect real-world legal intelligence Gursoy and Cai (2025).
+These MCQs assess multi-dimensional competencies cambria2024xai; corfmat2025high relevant to legal intelligence.
+Furthermore, to address cognitive coverage limitations in existing benchmarks, we propose a framework structured around the dimensions of legal theory and practice, organizing tasks and abilities to reflect real-world legal intelligence gursoy2025artificial.
 Focusing on Chinese laws ensures a robust and meaningful assessment, as distinct legal systems would otherwise dilute the evaluation.
-Leveraging LexGenius, we evaluated 12 state-of-the-art (SOTA) LLMs and 2 prompting strategies, including naive and chain-of-thought (CoT) prompting Kojima et al.
-(2022).
+Leveraging LexGenius, we evaluated 12 state-of-the-art (SOTA) LLMs and 2 prompting strategies, including naive and chain-of-thought (CoT) prompting kojima2022large.
 A baseline performance, constructed by 6 legal professionals, was also established for comparison.
-Results show that even the top-performing LLM, DeepSeek-R1 Bi et al.
-(2024), exhibits a significant gap compared to human experts across various legal general intelligence abilities Yao et al.
-(2025b); Hannah et al.
-(2025); Dong et al.
-(2025), as shown in Figure 1.
+Results show that even the top-performing LLM, DeepSeek-R1 bi2024deepseek, exhibits a significant gap compared to human experts across various legal general intelligence abilities yao2025intelligent; hannah2025legal; dong2025safeguarding, as shown in Figure 1.
 In summary, our contributions in this work include:
@@
 Recently, a series of legal benchmarks have emerged (see Table 1).
-They have made significant contributions to evaluating LLMs’ performance Kanapala et al.
-(2019); Yao et al.
-(2025a), including retrieval (STARD, LeCaRD) Su et al.
-(2024); Li et al.
-(2024c), question answering (JEC-QA, Legal CQA) Zhong et al.
-(2020); Askari et al.
-(2022), classification (LexGLUE) Chalkidis et al.
-(2021), reasoning (LegalBench, LexEval) Guha et al.
-(2023); Li et al.
-(2024a), and others (Laiw and LawBench) Dai et al.
-(2025); Fei et al.
-(2024).
-However, most benchmarks remain task-oriented and outcome-focused, offering limited insight into the underlying legal general intelligence of LLMs  Yue et al.
-(2024b).
+They have made significant contributions to evaluating LLMs’ performance kanapala2019text; yao2025elevating, including retrieval (STARD, LeCaRD) su2024stard; li2024lecardv2, question answering (JEC-QA, Legal CQA) zhong2020jec; askari2022expert, classification (LexGLUE) chalkidis2021lexglue, reasoning (LegalBench, LexEval) guha2023legalbench; li2024lexeval, and others (Laiw and LawBench) dai2025laiw; fei2024lawbench.
+However, most benchmarks remain task-oriented and outcome-focused, offering limited insight into the underlying legal general intelligence of LLMs  yue2024lawllm.
 Expert-level Benchmarks.
-To usher in the second half of AI, a series of expert-level benchmarks for evaluating LLMs have emerged across various domains Cao et al.
-(2025); Ni et al.
-(2025); Li et al.
-(2025c, d): PhysBench Chow et al.
-(2025) and PhysReason Zhang et al.
-(2025c) enhance LLMs’ understanding of physics; MedXpertQA Zuo et al.
-(2025) and MedagentsbenchTang et al.
-(2025) focus on medical knowledge; UGMathBench Xu et al.
-(2025b) assesses math reasoning; and UniToMBench Thiyagarajan et al.
-(2025) improves theory of mind.
-Benchmarks like ShotBench Liu et al.
-(2025a), FinTMMBench Zhu et al.
-(2025), and Chengyu-Bench Fu et al.
-(2025) evaluate other fields.
+To usher in the second half of AI, a series of expert-level benchmarks for evaluating LLMs have emerged across various domains cao2025toward; ni2025survey; li2025legalagentbench; li2025fundamental: PhysBench chow2025physbench and PhysReason zhang2025physreason enhance LLMs’ understanding of physics; MedXpertQA zuo2025medxpertqa and Medagentsbenchtang2025medagentsbench focus on medical knowledge; UGMathBench xu2025ugmathbench assesses math reasoning; and UniToMBench thiyagarajan2025unitombench improves theory of mind.
+Benchmarks like ShotBench liu2025shotbench, FinTMMBench zhu2025fintmmbench, and Chengyu-Bench fu2025chengyu evaluate other fields.
 However, an expert-level benchmark for legal intelligence is absent.
-Wang et al.
-(2025).
+wang2025survey.
 3 LexGenius Framework
 In this section, we outline the LexGenius framework (including seven dimensions, eleven tasks, and twenty abilities), which is shown in Figure 2.
-3.1 Dimension: Education and Career Drive
-The seven legal dimensions of LexGenius are based on Bloom’s Taxonomy of Educational Objectives Bloom et al.
-(1956), covering the cognitive hierarchy of remembering, understanding, applying, analyzing, evaluating, and creating, alongside the modular model used in legal evaluations across countries, focusing on normative understanding, rule application, procedural operation, and value judgment Wu and Chan (2012); Moon (2020); Parsons et al.
-(2024).
+3.1 Dimension: Education and Career Focus
+The seven legal dimensions of LexGenius are based on Bloom’s Taxonomy of Educational Objectives bloom1956taxonomy, covering the cognitive hierarchy of remembering, understanding, applying, analyzing, evaluating, and creating, alongside the modular model used in legal evaluations across countries, focusing on normative understanding, rule application, procedural operation, and value judgment wu2012regulatory; moon2020case; parsons2024georgia.
 In the hierarchy, remembering and understanding correspond to legal understanding, applying to legal application, analyzing to legal reasoning, evaluating to legal ethics and law and society, creating to advanced arguments, legal language to clarity, and judicial practice to procedural integrity, forming a framework aligned with cognitive principles and professional needs.
@@
 It includes three steps: data collection and structuring, construction of MCQs, and manual review.
-3.2 Task: Theory and Practice Coexist
-Further, based on Legal Hermeneutics Leyh (2021) and Problem-Solving Cycle theory Stein (1993), we decompose LexGenius’s 7 legal intelligence dimensions into 11 tasks.
+3.2 Task: Theory and Practice Synergy
+Further, based on Legal Hermeneutics leyh2021legal and Problem-Solving Cycle theory stein1993ideal, we decompose LexGenius’s 7 legal intelligence dimensions into 11 tasks.
 These tasks align with common legal practice requirements and focus on textual deconstruction, case adaptation, and procedural implementation.
@@
 3.3 Ability: Constructivist Learning-based
-Furthermore, based on Constructivist Learning Theory Ariati et al.
-(2025), we extract twenty atomic legal intelligence abilities from the eleven tasks.
+Furthermore, based on Constructivist Learning Theory ariati2025constructivist, we extract twenty atomic legal intelligence abilities from the eleven tasks.
 The theory shifts from outcome assessment to capturing knowledge paths through cognitive traces, simulating real legal scenarios to ensure that the evaluation reflects true professional abilities while aligning cognitive principles with occupational needs.
@@
 We avoided reusing existing legal datasets to minimize contamination risks and ensure originality.
-A structured legal competency framework, developed by experienced legal experts, was used to comprehensively cover core legal intelligence abilities Li et al.
-(2025f).
-To ensure benchmark effectiveness, a structured review process was established Mohammadi et al.
-(2025); Li et al.
-(2025b).
+A structured legal competency framework, developed by experienced legal experts, was used to comprehensively cover core legal intelligence abilities li2025survey.
+To ensure benchmark effectiveness, a structured review process was established mohammadi2025evaluation; li2025generation.
 Reviewers are master’s candidates in law, systematically trained and thoroughly familiar with key regulations, case frameworks, and legal reasoning methods.
@@
 Step 2: Construction of MCQs.
-There are two methods for constructing multiple-choice questions, both based on LLM.
+There are two methods for constructing multiple-choice questions, both based on the large language model.
 One is to screen and modify legal examination questions.
-Questions meeting the defined legal abilities are retained; others are modified using LLM.
-For questions related to legal soft intelligence Schiff et al.
-(2025), LLM is used to generate them.
+Questions meeting the legal abilities are retained; others are modified using LLM.
+For questions related to legal soft intelligence, an LLM is used to generate them.
 The first method selects and modifies questions from the legal question bank: MCQs with a single correct answer are retained.
 In contrast, those with multiple answers are adapted using LLMs and prompt templates to ensure fairness and difficulty.
-For LLM-generated MCQs, we designed prompt templates with task constraints, ability descriptions, and examples to guide question generation based on legal cases, ensuring unique answers and a clear legal basis.
+For LLM-generated MCQs, we designed prompt templates with task constraints, ability descriptions, and examples to guide question generation based on specific legal cases, ensuring unique answers and a clear legal basis.
 Step 3: Manual Review.
@@
 4.4 Evaluation Method
-To evaluate the LLM’s legal intelligence, LexGenius adopts a three-level framework Kahan (2015); Huang et al.
-(2024).
+To evaluate the LLM’s legal intelligence, LexGenius adopts a three-level framework kahan2015laws; huang2024cmdl.
 The dimension level categorizes tasks into legal cognition areas, providing insight into performance.
@@
 5.1 Experimental Setup
-We evaluated twelve SOTA LLMs with LexGenius, which include DeepSeek-LLM-7B-Chat (DeepSeek-7B) Bi et al.
-(2024), Qwen-2.5-7B-Instruct (Qwen-2.5-7B) Hui et al.
-(2024), Qwen-2.5-1.5B-Instruct (Qwen-2.5-1.5B) Hui et al.
-(2024), Qwen-3-8B Yang et al.
-(2025), Qwen-3-4B Yang et al.
-(2025), GLM-4-9B-Chat (GLM-4-9B) GLM et al.
-(2024), LLaMA-3.2-1B-Instruct (LLaMA-3.2-1B) Grattafiori et al.
-(2024), LLaMA-3.2-8B-Instruct (LLaMA-3.2-8B) Grattafiori et al.
-(2024), DeepSeek-R1 Guo et al.
-(2025), DeepSeek-V3 Liu et al.
-(2024), GPT-4o mini Hurst et al.
-(2024), and GPT‑4.1 nano Brown et al.
-(2020).
+We evaluated twelve SOTA LLMs with LexGenius, which include DeepSeek-LLM-7B-Chat (DeepSeek-7B) bi2024deepseek, Qwen-2.5-7B-Instruct (Qwen-2.5-7B) hui2024qwen2, Qwen-2.5-1.5B-Instruct (Qwen-2.5-1.5B) hui2024qwen2, Qwen-3-8B yang2025qwen3, Qwen-3-4B yang2025qwen3, GLM-4-9B-Chat (GLM-4-9B) glm2024chatglm, LLaMA-3.2-1B-Instruct (LLaMA-3.2-1B) grattafiori2024llama, LLaMA-3.2-8B-Instruct (LLaMA-3.2-8B) grattafiori2024llama, DeepSeek-R1 guo2025deepseek, DeepSeek-V3 liu2024deepseek, GPT-4o mini hurst2024gpt, and GPT‑4.1 nano brown2020language.
 We followed the official protocols, using official APIs or LLM weights where applicable.
@@
 5.5 With Different Enhanced Methods (RQ4)
-As shown in Table 5, the comparison results (more details are in Appendix E) reveal a triple decoupling phenomenon in LLM legal intelligence: (i) Scale-performance decoupling shows that the Qwen2.5 series outperforms the Qwen3 series, with legal intelligence exhibiting a non-linear relationship with model size, indicating that domain-specific pretraining, not scale, drives legal abilities.
+As shown in Table 5, the comparison results (more details are in Appendix F) reveal a triple decoupling phenomenon in LLM legal intelligence: (i) Scale-performance decoupling shows that the Qwen2.5 series outperforms the Qwen3 series, with legal intelligence exhibiting a non-linear relationship with model size, indicating that domain-specific pretraining, not scale, drives legal abilities.
 (ii) Reasoning paradigm decoupling shows that CoT leads to negative transfer in legal tasks due to deterministic constraints, causing probabilistic exploration and semantic drift in closed-solution tasks.
@@
 Experimental results reveal significant legal intelligence gaps of LLMs, highlighting disparities with human legal experts and their specific weaknesses in legal general intelligence.
+Limitations
+In Appendix H, we discuss the limitations of LexGenius. Furthermore, we outline the future work of LexGenius in Appendix I.
+Ethical Considerations
+This work complies with the ACL Ethics Policy, relying on anonymized, publicly available legal resources to ensure privacy and academic integrity.
 Appendix
@@
 Legal general intelligence is not a stack of tasks, but a simulation of cognitive collaborative chains.
-Existing evaluation frameworks for legal intelligence often operate at the level of classification or question-answering tasks, presenting only the model’s macro-level accuracy on isolated benchmarks.
-Such metrics fail to uncover where a model goes wrong within the complex cognitive chain of legal decision-making.
-Legal intelligence is not a mere sum of discrete tasks—it is an organic coordination of systematic cognitive capabilities, spanning multiple stages such as statutory interpretation, fact extraction, rule adaptation, ethical judgment, and social impact assessment.
-The performance of an LLM should therefore be decomposed into a multi-stage flow of abilities, rather than treated as a monolithic output.
-Accordingly, an effective evaluation framework must faithfully map this cognitive chain, avoiding the pitfall of reducing legal intelligence to mere task-solving while overlooking how the model actually thinks.
+Existing frameworks often focus on classification or question-answering tasks, presenting only macro-level accuracy on isolated benchmarks.
+Such metrics fail to pinpoint errors within the complex cognitive chain of legal decision-making.
+Legal intelligence is not a sum of discrete tasks but an organic coordination of systematic capabilities, spanning statutory interpretation, fact extraction, rule adaptation, ethical judgment, and social impact assessment.
+LLM performance should therefore be decomposed into a multi-stage flow, rather than treated as monolithic.
+Accordingly, effective evaluation must map this cognitive chain, avoiding the pitfall of reducing legal intelligence to task-solving while ignoring how the model thinks.
 From performance reporting to capability diagnosis and explanation.
-Traditional metrics such as accuracy or F1 score can indicate whether a model’s output is right or wrong, but they fail to address a more fundamental question: At which cognitive stage did the model fail?
+Traditional metrics like accuracy or F1 indicate output correctness, but they fail to address a fundamental question: At which cognitive stage did the model fail?
 Was it a semantic misunderstanding?
-A misapplication of legal rules?
+Rule misapplication?
 Or an ethical blind spot?
 LexGenius introduces 20 atomic legal general intelligence abilities and establishes an interpretable, traceable, and auditable evaluation system through a tri-level mapping mechanism across the ability layer, task layer, and cognitive dimension.
-This structure not only reveals the model’s deficiencies in specific micro-level legal abilities but also provides an actionable feedback loop for subsequent capability-oriented training, prompt optimization, and safety enhancement.
+This structure reveals deficiencies in specific micro-level abilities and provides an actionable feedback loop for capability-oriented training, prompt optimization, and safety enhancement.
 Enabling cross-stage cognitive analysis and transferability research.
 The complexity of legal reasoning lies in its chained structure: statutory semantics $\rightarrow$ case fact abstraction $\rightarrow$ rule application $\rightarrow$ ethical judgment $\rightarrow$ precedent alignment.
-An evaluation framework that fails to distinguish performance across these stages cannot effectively support research in multi-hop reasoning, chain-of-thought attention, or multi-task learning.
-Through hierarchical abstraction and modular decomposition, our framework standardizes this cognitive pathway, offering a clear experimental baseline for subsequent investigations into how models learn to transfer knowledge and whether they exhibit generalizable reasoning capabilities.
+Frameworks failing to distinguish performance across these stages cannot support research in multi-hop reasoning, chain-of-thought attention, or multi-task learning.
+Through hierarchical abstraction and modular decomposition, our framework standardizes this cognitive pathway, offering a clear baseline for investigating knowledge transfer and generalizable reasoning capabilities.
 Toward a professional-grade legal general intelligence evaluation paradigm.
-As LLMs begin to approach the professional thresholds of bar examinations and real-world legal practice, evaluation frameworks must likewise advance to a professional-grade level.
-LexGenius draws on standards from the National Legal Professional Qualification Examination and international bar exams, distinguishing among multiple specialized dimensions of legal competence, such as legal semantic understanding, norm alignment, ethical judgment, and procedural compliance.
-This framework breaks away from the general-purpose perspective of traditional NLP benchmarks, aligning instead with the practical demands of legal work.
-Such a professional, ability-oriented, structured evaluation not only reveals the current boundaries of model capabilities but also illuminates their potential trajectories of future development.
+As LLMs approach the professional thresholds of bar examinations and real-world legal practice, evaluation frameworks must likewise advance to a professional-grade level.
+LexGenius draws on standards from the National Legal Professional Qualification Examination and international bar exams, distinguishing specialized dimensions of legal competence, such as legal semantic understanding, norm alignment, ethical judgment, and procedural compliance.
+This framework breaks from the general-purpose perspective of traditional NLP benchmarks, aligning with the practical demands of legal work.
+Such a professional, ability-oriented, structured evaluation reveals current model boundaries and illuminates potential development trajectories.
+Level Description Goals and Functions Number Dimension Core legal intelligence focus areas and judgments (e.g., legal reasoning) Build top-level cognition and capability aggregation 7 Task Scenario-based legal tasks (e.g., case reasoning and judgment prediction) Mid-level structure linking abilities and test design 11 Ability Measurable legal intelligence ability units (e.g., statute understanding and interpretation) Minimum unit, supporting fine-grained evaluation and diagnosis 20
+Table 6: Hierarchical levels of the LexGenius and corresponding implementation counts.
+It includes the Dimension level (high-level cognitive targets), the Task level (scenario-based applications), and the Ability level (fine-grained evaluable units), along with the number of implemented benchmarks under each category.
 A.2 Framework hierarchy and implementation
@@
 Appendix B Definitions of Legal Intelligence Abilities
-This section provides detailed definitions for the 20 atomic legal intelligence abilities in the LexGenius framework, following the ability names used in Figure 2.
+This section provides detailed definitions for the twenty atomic legal intelligence abilities in the proposed LexGenius framework, following the ability names utilized in Figure 2.
 Each ability represents a measurable unit of legal general intelligence, assessed through standardized multiple-choice questions.
@@
 Figure 28: The MCQ sample of ability 20. The left is the original text, and the right is the English translation.
-Appendix C Experimental setup details
-C.1 Large Language Models
-We evaluated twelve popular and SOTA LLMs on the proposed LexGenius.
-For GPT-4o mini, GPT-4 nano, DeepSeek V3, and DeepSeek R1, we accessed the models via their official APIs.
+Appendix C Annotation Details
+We recruited nine master’s candidates in law for double-blind annotation.
+As detailed in Section 4.2 and Figure 3, evaluators assessed questions on a 5-point scale based on five specific criteria: statutory accuracy, logical soundness, answer uniqueness, competence alignment, and normative wording.
+Arbitration was triggered for score discrepancies exceeding two points.
+Participants were informed of data usage and management policies.
+To ensure fair labor practices, annotators were compensated at a rate of US\$15 per hour, which exceeds the local average wage for student research assistants.
+Appendix D Experimental setup details
+D.1 Large Language Models
+We evaluated 12 LLMs on LexGenius.
+For GPT-4o mini, GPT-4 nano, DeepSeek-V3, and DeepSeek-R1, we accessed them via their official APIs.
 For other LLMs, we conducted experimental tests using the official weights.
-The twelve popular SOTA LLMs are as follows:
-Qwen2.5-1.5B-Instruct Hui et al.
-(2024): A lightweight instruction-tuned model released by Alibaba with 1.5B parameters, designed for edge deployment and local inference with bilingual support and basic task execution.
-Qwen2.5-7B-Instruct Hui et al.
-(2024): A mid-scale model in the Qwen2.5 series, optimized for stronger reasoning and instruction following, suitable for more complex language tasks in medium-sized deployments.
-Qwen3-4B Yang et al. (2025): A 4B-parameter model from the third-generation Qwen series, showing strong performance in multilingual, coding, and logical tasks.
-Qwen3-8B Yang et al. (2025): An enhanced version of Qwen3 with extended context length and multilingual capabilities, significantly improving performance in complex reasoning and generation tasks.
-LLaMA-3.2-1B-Instruct Grattafiori et al.
-(2024): A compact instruction-tuned model from Meta’s LLaMA 3 series, designed for resource-constrained environments while maintaining core instruction-following capabilities.
-LLaMA-3.2-8B-Instruct Grattafiori et al.
-(2024): A standard model in the LLaMA 3 lineup, offering high-quality multilingual understanding, code generation, and reasoning, achieving state-of-the-art results across many tasks.
-GLM-4-9B-Chat GLM et al.
-(2024): A 9B bilingual chat model developed by Zhipu AI, equipped for multi-turn dialogue, tool use, and contextual memory, with strong performance particularly in Chinese semantic understanding.
-DeepSeek-LLM-7B-Chat Bi et al.
-(2024): A 7B bilingual chat model by DeepSeek, integrating capabilities in code generation, mathematics, and language understanding, suitable for dialogue and multitask settings.
-DeepSeek-R1 Guo et al.
-(2025): A large-scale open-source language model developed by the Chinese company DeepSeek, featuring strong capabilities in mathematics, programming, and reasoning with efficient training and leading performance.
-DeepSeek-V3 Liu et al.
-(2024): A large language model based on a mixture-of-experts architecture, excelling in mathematics, programming, and logical reasoning, and is well-suited for a variety of intelligent application scenarios.
-GPT-4o mini Hurst et al.
-(2024): A parameter-efficient version of OpenAI’s GPT-4o model, supporting multimodal inputs (text, image, audio) with consistent alignment behavior as its full-size counterpart.
-GPT-4.1 nano Hurst et al.
-(2024): An ultra-compact model in the GPT-4.1 series, designed for on-device and embedded inference with support for moderate-length contexts and basic reasoning under resource constraints.
+These LLMs are as follows:
+Qwen2.5-1.5B-Instruct hui2024qwen2: A lightweight instruction-tuned model released by Alibaba with 1.5B parameters, designed for edge deployment and local inference with bilingual support and basic task execution.
+Qwen2.5-7B-Instruct hui2024qwen2: A mid-scale model in the Qwen2.5 series, optimized for stronger reasoning and instruction following, suitable for more complex language tasks in medium-sized deployments.
+Qwen3-4B yang2025qwen3: A 4B-parameter model from the third-generation Qwen series, showing strong performance in multilingual, coding, and logical tasks.
+Qwen3-8B yang2025qwen3: An enhanced version of Qwen3 with extended context length and multilingual capabilities, significantly improving performance in reasoning and generation tasks.
+LLaMA-3.2-1B-Instruct grattafiori2024llama: A compact instruction-tuned model from Meta’s LLaMA 3 series, designed for resource-constrained environments while maintaining core instruction-following capabilities.
+LLaMA-3.2-8B-Instruct grattafiori2024llama: A standard model in the LLaMA 3 lineup, offering high-quality multilingual understanding, code generation, and reasoning, achieving state-of-the-art results across many tasks.
+GLM-4-9B-Chat glm2024chatglm: A 9B bilingual chat model developed by Zhipu AI, equipped for multi-turn dialogue, tool use, and contextual memory, with strong performance particularly in Chinese semantic understanding.
+DeepSeek-LLM-7B-Chat bi2024deepseek: A 7B bilingual chat model by DeepSeek, integrating capabilities in code generation, mathematics, and language understanding, suitable for dialogue and multitask settings.
+DeepSeek-R1 guo2025deepseek: A large-scale open-source language model developed by the Chinese company DeepSeek, featuring strong capabilities in mathematics, programming, and reasoning with efficient training and leading performance.
+DeepSeek-V3 liu2024deepseek: A large language model based on a mixture-of-experts architecture, excelling in mathematics, programming, and logical reasoning, and is well-suited for a variety of intelligent application scenarios.
+GPT-4o mini hurst2024gpt: A parameter-efficient version of OpenAI’s GPT-4o, supporting multimodal inputs (text, image, audio) with consistent alignment behavior as its full-size counterpart.
+GPT-4.1 nano hurst2024gpt: An ultra-compact model in the GPT-4.1 series, designed for on-device and embedded inference with support for moderate-length contexts and basic reasoning under resource constraints.
 Figure 29: The two utilized prompt methods for LLMs. In this figure, we provide the Chinese and English texts.
-C.2 Two prompt methods
-We employed two prompting strategies to evaluate the legal general intelligence on the LexGenius.
-One is a simple prompt (naive), and the other is a Chain-of-Thought (CoT) prompt.
-These two strategies are illustrated in Figure 29.
-Naive Prompt is the most basic form of prompt design, where the task objective is directly stated to the language model, with the expectation that it will generate the final output without revealing any intermediate reasoning steps.
-This approach relies on the model’s end-to-end learning ability, requiring it to complete the full input-to-output mapping within a single forward pass.
-Its main advantage lies in simplicity and speed, but it places high demands on the model’s implicit knowledge and generalization capability.
-Especially for complex tasks, it is prone to errors due to semantic ambiguity or interference from multi-step reasoning chains.
-Conceptually, this prompting strategy does not explicitly activate the model’s thinking path, which may lead it to overlook latent contextual connections, implicit assumptions, or task structure.
-As task complexity increases, performance tends to degrade rapidly due to the lack of a mechanism for step-by-step reasoning construction.
-CoT Prompt is to explicitly guide the model through step-by-step reasoning by decomposing a task into a sequence of intermediate steps expressed in natural language Kojima et al.
-(2022).
-This approach encourages the model to simulate the human problem-solving process and functions as a form of prompt programming, activating internal causal chains and logical pathways by embedding intermediate reasoning states within the prompt.
-Mechanistically, CoT prompting leverages the structured knowledge and implicit logical modeling capabilities that large models naturally acquire during training.
-When the prompt includes cues such as think, step, or other structured instructions, the model tends to produce coherent reasoning processes rather than skipping directly to the final answer.
-This reasoning chain helps reduce error rates in complex tasks and improves the model’s ability to handle multi-step dependencies, conditional logic, and task decomposition.
-Appendix D Results of 20 Legal Intelligence Abilities
+Ability Human LLM 1 LLM 2 LLM 3 LLM 4 LLM 5 LLM 6 LLM 7 LLM 8 LLM 9 LLM 10 LLM 11 LLM 12
+Naive Prompt
+Ability 1 89.63 45.00 58.20 50.80 55.20 28.80 37.80 53.00 39.00 64.40 64.80 47.00 41.40
+Ability 2 87.41 37.78 60.89 46.22 54.22 30.67 33.33 46.22 34.22 67.11 65.78 47.11 40.44
+Ability 3 90.37 49.38 65.15 62.66 65.15 28.63 39.83 59.34 38.59 70.54 70.54 55.19 57.68
+Ability 4 92.59 39.40 56.60 48.00 50.60 25.80 32.20 47.60 32.20 59.20 58.00 46.00 39.00
+Ability 5 88.89 43.45 55.36 52.38 50.60 23.81 27.98 41.67 35.71 60.71 60.12 39.29 47.62
+Ability 6 85.93 40.19 60.29 50.72 55.02 22.49 33.97 54.07 34.93 63.64 64.59 44.50 43.06
+Ability 7 85.93 36.20 47.80 41.40 42.80 25.60 31.60 39.40 35.80 46.80 46.80 40.40 35.40
+Ability 8 92.59 39.60 52.20 46.00 48.40 23.20 34.80 42.60 27.40 53.00 55.20 40.40 35.80
+Ability 9 88.89 31.03 35.78 29.74 25.00 30.60 33.19 24.57 25.43 37.93 37.07 30.17 21.55
+Ability 10 86.67 59.00 64.80 58.00 64.20 46.80 56.20 66.00 48.40 67.60 67.60 65.20 59.20
+Ability 11 83.70 52.40 63.20 63.80 61.60 42.00 50.80 56.80 41.80 74.00 73.00 59.20 54.40
+Ability 12 89.63 43.55 63.55 50.65 60.97 27.74 29.35 58.71 32.26 74.19 73.55 45.81 45.81
+Ability 13 85.93 60.40 68.60 64.60 65.20 51.00 63.40 66.20 46.80 68.00 67.40 69.20 60.40
+Ability 14 87.41 57.20 71.00 69.00 67.60 45.80 55.00 64.40 51.60 76.80 76.40 67.60 63.00
+Ability 15 91.85 56.40 58.00 55.60 54.80 46.00 51.60 57.60 40.20 65.40 66.20 61.20 54.40
+Ability 16 82.96 64.40 69.20 62.00 61.20 59.00 59.60 63.60 49.20 75.00 72.60 67.60 59.00
+Ability 17 89.63 38.40 45.80 40.80 45.00 23.80 34.20 39.40 33.40 48.80 49.40 40.20 37.00
+Ability 18 90.37 50.80 64.60 58.20 66.00 44.00 53.60 60.00 40.40 72.40 71.80 60.40 59.00
+Ability 19 79.26 58.20 74.00 65.80 69.60 37.80 54.60 63.80 45.60 80.00 81.20 65.80 61.80
+Ability 20 87.41 34.60 45.80 41.40 40.80 26.60 28.80 39.00 29.40 48.40 47.20 42.60 35.00
+CoT Prompt
+Ability 1 89.63 45.00 57.60 50.60 54.40 30.00 36.40 53.20 36.80 64.40 64.20 48.00 56.80
+Ability 2 87.41 38.67 60.00 44.44 55.11 28.44 32.00 45.78 30.67 66.67 66.67 46.22 50.22
+Ability 3 90.37 48.13 63.90 62.24 66.80 26.56 38.59 60.17 33.20 72.20 71.78 55.19 56.85
+Ability 4 92.59 42.00 56.20 48.00 49.20 23.60 32.20 47.00 32.00 60.40 58.80 45.80 40.00
+Ability 5 88.89 42.26 54.17 52.98 50.60 28.57 30.95 41.07 27.38 59.52 61.31 40.48 47.62
+Ability 6 85.93 42.58 59.33 50.72 58.37 25.36 35.41 54.55 33.49 64.59 64.59 45.45 44.02
+Ability 7 85.93 36.00 48.00 42.00 41.40 24.60 32.00 39.60 34.40 46.60 47.60 39.40 39.40
+Ability 8 92.59 39.60 51.80 44.20 47.40 22.20 35.00 41.80 29.80 53.60 53.40 40.60 48.00
+Ability 9 88.89 31.90 35.78 28.02 22.84 31.90 33.19 28.02 28.45 37.50 38.79 29.31 32.76
+Ability 10 86.67 59.40 65.40 58.60 62.80 46.60 55.20 65.60 50.20 68.00 67.40 65.80 60.80
+Ability 11 83.70 52.60 61.80 63.40 60.40 40.20 51.00 57.20 36.40 73.20 71.40 58.60 58.40
+Ability 12 89.63 44.84 63.23 50.32 60.65 26.45 29.03 59.68 32.58 74.19 73.23 46.45 60.97
+Ability 13 85.93 61.80 68.60 64.40 63.20 53.40 64.20 66.80 49.00 68.20 67.80 68.60 64.40
+Ability 14 87.41 57.20 70.40 68.20 66.00 41.40 56.40 63.40 50.40 77.00 77.20 67.20 66.20
+Ability 15 91.85 56.20 59.60 54.00 54.60 45.60 51.80 57.40 42.60 65.00 65.20 61.40 54.00
+Ability 16 82.96 66.00 68.80 61.20 60.80 56.80 61.20 63.60 49.20 75.00 75.40 68.80 61.40
+Ability 17 89.63 38.80 45.20 41.20 44.40 21.60 32.40 38.40 30.60 51.20 48.80 39.60 37.00
+Ability 18 90.37 50.80 64.60 58.00 65.00 42.80 54.80 60.80 42.60 73.60 72.80 61.80 60.60
+Ability 19 79.26 59.20 74.40 66.00 68.40 39.00 56.60 64.60 43.60 79.60 81.00 67.00 66.00
+Ability 20 87.41 36.60 46.20 41.80 39.80 25.60 30.40 40.40 29.40 47.00 47.40 42.00 46.80
+Table 7: Comparison of performance across 20 legal intelligence abilities for Naive Prompt and CoT Prompt on various LLMs (all values in %).
+LLM 1 is Qwen2.5-1.5B-Instruct; LLM 2 is Qwen2.5-7B-Instruct; LLM 3 is Qwen3-4B; LLM 4 is Qwen3-8B; LLM 5 is Llama-3.2-1B-Instruct; LLM 6 is Llama-3.2-8B-Instruct; LLM 7 is GLM-4-9B-Chat; LLM 8 is DeepSeek-LLM-7B-Chat; LLM 9 is DeepSeek-R1; LLM 10 is DeepSeek-V3; LLM 11 is GPT-4o mini; and LLM 12 is GPT-4.1 nano.
+D.2 Two Prompt Methods
+To evaluate LexGenius, we employed Naive and CoT strategies (see Figure 29).
+The Naive prompt prioritizes efficiency via direct output generated in a single pass, though its lack of reasoning often degrades performance on complex tasks.
+Conversely, CoT simulates human problem-solving by decomposing tasks into intermediate steps, activating causal chains to significantly reduce errors in multi-step dependencies and conditional logic.
+Appendix E Results of Twenty Abilities
 Based on the structured capability framework provided by LexGenius, we evaluated the performance of various SOTA LLMs across 20 legal intelligence abilities (see Table 7).
@@
 These dimensions are further divided into 11 tasks and 20 abilities, covering a comprehensive range of legal intelligence abilities, including understanding, reasoning, application, ethical judgment, language processing, socio-legal interaction, and judicial practice.
-The results of 20 legal intelligence abilities for the 12 SOTA LLMs are as illustrated in Table 7.
+The results of 20 legal intelligence abilities for the 12 LLMs are shown in Table 7.
 The LLMs’ legal intelligence abilities decline significantly in tasks requiring deeper abstraction.
@@
 This highlights the need for further model optimization in sociological, ethical, and institutional aspects of legal general intelligence.
-Level Description Goals and Functions Number Dimension Core legal intelligence focus areas and judgments (e.g., legal reasoning) Build top-level cognition and capability aggregation 7 Task Scenario-based legal tasks (e.g., case reasoning and judgment prediction) Mid-level structure linking abilities and test design 11 Ability Measurable legal intelligence ability units (e.g., statute understanding and interpretation) Minimum unit, supporting fine-grained evaluation and diagnosis 20
-Table 6: Hierarchical levels of the LexGenius and corresponding implementation counts.
-It includes the Dimension level (high-level cognitive targets), the Task level (scenario-based applications), and the Ability level (fine-grained evaluable units), along with the number of implemented benchmarks under each category.
-Figure 30: The correlation analysis of legal intelligence ability, task, and dimension in LexGenius for 12 LLMs.
-Appendix E With Different Enhanced Methods
-To accurately compare and evaluate the impact of different optimization and enhancement methods on the legal intelligence capabilities of Large Language Models (LLMs), we selected four LLMs (including Qwen2.5-1.5B-Instruct, Qwen2.5-7B-Instruct, Qwen3-4B, and Qwen3-8B) and experimented with Supervised Fine-Tuning (SFT), Chain-of-Thought (CoT), Retrieval-Augmented Generation (RAG), and Reinforcement Learning (RL) algorithms.
-We randomly sampled 64 test instances from each of the 20 ability test sets in LexGenius, resulting in 1,280 total samples for evaluation.
-The remaining 7,105 data samples were used as the training set for SFT and RL, as well as for constructing the retrieval corpus.
-The experimental results of these LLMs, after applying these enhancement methods across various dimensions and tasks of legal intelligence, are illustrated in Table 8 and Table 9.
-Appendix F Correlation Analysis
-The average performance of the twelve SOTA LLMs on LexGenius is utilized to conduct correlation analysis on the benchmark (as shown in Figure 30).
-It illustrates that most of the legal intelligence abilities (left), tasks (upper right), and dimensions (lower right) exhibit low correlations.
-It shows the effectiveness of the developed questions in our LexGenius, because the low intercorrelation suggests that models cannot rely on general legal heuristics or shallow transfer across domains to perform well; instead, success in one category does not guarantee success in others.
-This reflects the comprehensive coverage and conceptual independence of our benchmark dimensions, further validating their robustness as an evaluation framework.
-Appendix G Limitation and Future Work
-To bridge the gap between current benchmark capabilities and the demands of real-world legal practice, we outline several directions for future work.
-While LexGenius establishes a foundational framework for evaluating legal general intelligence, it remains constrained by unimodal design, limited jurisdictional scope, and a lack of temporal modeling.
-To address these limitations, future iterations will pursue enhancements in three key areas: incorporating multimodal legal tasks, expanding linguistic and legal diversity, and introducing dynamic temporal reasoning.
-These improvements aim to more accurately reflect the complexity of legal environments and to enable more robust, generalizable evaluations of large language models.
-We hope that LexGenius and these discussions can promote the development and research of legal general intelligence.
-G.1 Limitation
-While LexGenius offers a structured framework for evaluating legal general intelligence in large language models, it still falls short in key areas that limit its realism and generalizability.
-Specifically, it lacks support for multimodal inputs, is confined to a single language and legal system, and does not account for the temporal dynamics of law.
-These gaps constrain its alignment with real-world legal complexity and hinder its ability to fully assess models’ applicability in diverse, evolving legal contexts.
-The following subsections detail these core limitations.
-Lack of Multimodal Tasks Limits Realistic Evidence Modeling.
-The current version of LexGenius is built and evaluated entirely on pure textual materials, without incorporating multimodal evidence types commonly found in real-world cases, such as scanned contract images, video stills, or audio transcriptions.
-This unimodal design fails to assess a model’s capabilities in visual perception, auditory understanding, and cross-modal reasoning, which are essential for handling actual judicial cases.
-The absence of multimodal inputs limits the applicability of LLMs in tasks such as evidence review, fact reconstruction, and the interpretation of visual-legal content, thereby reducing the fidelity of the evaluation to real-world legal scenarios.
-Linguistic and Jurisdictional Limitations Undermine Cross-Cultural Generalization.
-LexGenius is currently constructed solely from Chinese-language corpora and based on the civil law system of Mainland China, exhibiting clear singularity in both linguistic and legal dimensions.
-As a result, model evaluations are only valid within the context of "Chinese civil law" and fail to capture broader capabilities such as interpreting international statutes, translating case law, or conducting comparative legal analysis.
-In tasks involving global legal services, cross-border disputes, or foreign compliance, this limitation significantly hampers a model’s transferability and generalization, impeding its evolution into a truly universal system of legal intelligence.
-Ability Human LLM 1 LLM 2 LLM 3 LLM 4 LLM 5 LLM 6 LLM 7 LLM 8 LLM 9 LLM 10 LLM 11 LLM 12 Naive Prompt Ability 1 89.63 45.00 58.20 50.80 55.20 28.80 37.80 53.00 39.00 64.40 64.80 47.00 41.40 Ability 2 87.41 37.78 60.89 46.22 54.22 30.67 33.33 46.22 34.22 67.11 65.78 47.11 40.44 Ability 3 90.37 49.38 65.15 62.66 65.15 28.63 39.83 59.34 38.59 70.54 70.54 55.19 57.68 Ability 4 92.59 39.40 56.60 48.00 50.60 25.80 32.20 47.60 32.20 59.20 58.00 46.00 39.00 Ability 5 88.89 43.45 55.36 52.38 50.60 23.81 27.98 41.67 35.71 60.71 60.12 39.29 47.62 Ability 6 85.93 40.19 60.29 50.72 55.02 22.49 33.97 54.07 34.93 63.64 64.59 44.50 43.06 Ability 7 85.93 36.20 47.80 41.40 42.80 25.60 31.60 39.40 35.80 46.80 46.80 40.40 35.40 Ability 8 92.59 39.60 52.20 46.00 48.40 23.20 34.80 42.60 27.40 53.00 55.20 40.40 35.80 Ability 9 88.89 31.03 35.78 29.74 25.00 30.60 33.19 24.57 25.43 37.93 37.07 30.17 21.55 Ability 10 86.67 59.00 64.80 58.00 64.20 46.80 56.20 66.00 48.40 67.60 67.60 65.20 59.20 Ability 11 83.70 52.40 63.20 63.80 61.60 42.00 50.80 56.80 41.80 74.00 73.00 59.20 54.40 Ability 12 89.63 43.55 63.55 50.65 60.97 27.74 29.35 58.71 32.26 74.19 73.55 45.81 45.81 Ability 13 85.93 60.40 68.60 64.60 65.20 51.00 63.40 66.20 46.80 68.00 67.40 69.20 60.40 Ability 14 87.41 57.20 71.00 69.00 67.60 45.80 55.00 64.40 51.60 76.80 76.40 67.60 63.00 Ability 15 91.85 56.40 58.00 55.60 54.80 46.00 51.60 57.60 40.20 65.40 66.20 61.20 54.40 Ability 16 82.96 64.40 69.20 62.00 61.20 59.00 59.60 63.60 49.20 75.00 72.60 67.60 59.00 Ability 17 89.63 38.40 45.80 40.80 45.00 23.80 34.20 39.40 33.40 48.80 49.40 40.20 37.00 Ability 18 90.37 50.80 64.60 58.20 66.00 44.00 53.60 60.00 40.40 72.40 71.80 60.40 59.00 Ability 19 79.26 58.20 74.00 65.80 69.60 37.80 54.60 63.80 45.60 80.00 81.20 65.80 61.80 Ability 20 87.41 34.60 45.80 41.40 40.80 26.60 28.80 39.00 29.40 48.40 47.20 42.60 35.00 CoT Prompt Ability 1 89.63 45.00 57.60 50.60 54.40 30.00 36.40 53.20 36.80 64.40 64.20 48.00 56.80 Ability 2 87.41 38.67 60.00 44.44 55.11 28.44 32.00 45.78 30.67 66.67 66.67 46.22 50.22 Ability 3 90.37 48.13 63.90 62.24 66.80 26.56 38.59 60.17 33.20 72.20 71.78 55.19 56.85 Ability 4 92.59 42.00 56.20 48.00 49.20 23.60 32.20 47.00 32.00 60.40 58.80 45.80 40.00 Ability 5 88.89 42.26 54.17 52.98 50.60 28.57 30.95 41.07 27.38 59.52 61.31 40.48 47.62 Ability 6 85.93 42.58 59.33 50.72 58.37 25.36 35.41 54.55 33.49 64.59 64.59 45.45 44.02 Ability 7 85.93 36.00 48.00 42.00 41.40 24.60 32.00 39.60 34.40 46.60 47.60 39.40 39.40 Ability 8 92.59 39.60 51.80 44.20 47.40 22.20 35.00 41.80 29.80 53.60 53.40 40.60 48.00 Ability 9 88.89 31.90 35.78 28.02 22.84 31.90 33.19 28.02 28.45 37.50 38.79 29.31 32.76 Ability 10 86.67 59.40 65.40 58.60 62.80 46.60 55.20 65.60 50.20 68.00 67.40 65.80 60.80 Ability 11 83.70 52.60 61.80 63.40 60.40 40.20 51.00 57.20 36.40 73.20 71.40 58.60 58.40 Ability 12 89.63 44.84 63.23 50.32 60.65 26.45 29.03 59.68 32.58 74.19 73.23 46.45 60.97 Ability 13 85.93 61.80 68.60 64.40 63.20 53.40 64.20 66.80 49.00 68.20 67.80 68.60 64.40 Ability 14 87.41 57.20 70.40 68.20 66.00 41.40 56.40 63.40 50.40 77.00 77.20 67.20 66.20 Ability 15 91.85 56.20 59.60 54.00 54.60 45.60 51.80 57.40 42.60 65.00 65.20 61.40 54.00 Ability 16 82.96 66.00 68.80 61.20 60.80 56.80 61.20 63.60 49.20 75.00 75.40 68.80 61.40 Ability 17 89.63 38.80 45.20 41.20 44.40 21.60 32.40 38.40 30.60 51.20 48.80 39.60 37.00 Ability 18 90.37 50.80 64.60 58.00 65.00 42.80 54.80 60.80 42.60 73.60 72.80 61.80 60.60 Ability 19 79.26 59.20 74.40 66.00 68.40 39.00 56.60 64.60 43.60 79.60 81.00 67.00 66.00 Ability 20 87.41 36.60 46.20 41.80 39.80 25.60 30.40 40.40 29.40 47.00 47.40 42.00 46.80
-Table 7: Comparison of performance across 20 legal intelligence abilities for Naive Prompt and CoT Prompt on various LLMs (all values in %).
-LLM 1 is Qwen2.5-1.5B-Instruct; LLM 2 is Qwen2.5-7B-Instruct; LLM 3 is Qwen3-4B; LLM 4 is Qwen3-8B; LLM 5 is Llama-3.2-1B-Instruct; LLM 6 is Llama-3.2-8B-Instruct; LLM 7 is GLM-4-9B-Chat; LLM 8 is DeepSeek-LLM-7B-Chat; LLM 9 is DeepSeek-R1; LLM 10 is DeepSeek-V3; LLM 11 is GPT-4o mini; and LLM 12 is GPT-4.1 nano.
-Lack of Evaluation on Temporal Sensitivity and Legal Validity Awareness.
-A core characteristic of law is its temporal nature—the applicable rules for a given issue may vary significantly across time, especially before and after legislative amendments.
-LexGenius currently does not incorporate a systematic temporal dimension to assess whether models can understand the time-bound applicability of statutes, the validity period of precedents, or transitional legal provisions.
-Without such temporal sensitivity tests, models may produce outdated or legally invalid answers when facing evolving legal frameworks, with no mechanism to detect these errors.
-G.2 Future work
-LexGenius aims to advance Chinese legal intelligence toward generalization.
-The current version covers fundamental legal domains and establishes a Dimension–Task–Ability evaluation framework, yet it still falls short of capturing the full complexity of real-world legal systems.
-Therefore, our future work will primarily focus on the following:
-Incorporating Multimodal Tasks to Enhance Realistic Evidence Modeling.
-The current version is constructed solely on textual materials and does not include multimodal information commonly encountered in real legal cases, such as scanned contracts, courtroom audio transcriptions, or surveillance video stills.
-The absence of such inputs limits the evaluation of model capabilities in visual perception, auditory comprehension, and cross-modal reasoning, all of which are essential for key legal tasks such as evidence review, fact reconstruction, and interpretation of visual-legal content.
-In future iterations, we plan to embed images, audio, and other modalities into tasks to assess models’ ability to reason and judge based on heterogeneous, multi-source information, thus aligning evaluation more closely with practical judicial needs.
-Expanding Linguistic and Jurisdictional Coverage to Improve Cross-Cultural Generalization.
-The current dataset is grounded in Chinese-language texts and the civil law system of Mainland China, exhibiting limitations in both language and legal tradition.
-This restricts the applicability of evaluation outcomes.
-Future versions will incorporate legal texts from Hong Kong, Macau, and Taiwan, as well as English-language statutes and case law from common law systems.
-We aim to construct bilingual legal QA pairs, statute translation tasks, and comparative law analyses to support the evaluation of models’ capabilities in understanding, aligning, and adapting across legal and linguistic contexts.
-This expansion will contribute to the training and benchmarking of legal language models equipped for global legal services.
-Introducing Dynamic Testing of Legal Temporality and Time Sensitivity.
-Legal applicability is highly time-dependent—amendments to laws can lead to drastically different rulings, and precedents often carry specific periods of validity and applicability.
-Currently, LexGenius lacks a systematic temporal dimension, making it difficult to evaluate whether a model can identify the applicable time windows of statutes, conditions for transitional provisions, or conflicts between old and new laws.
-Future versions will include temporally structured legal tasks that require models to make dynamic judgments under varying legal timeframes, thereby enhancing their understanding and adaptability to evolving legal systems.
 Model Legal Und. Legal Rea. Legal App. Legal Ethics Legal Lan. Law & Soc. Judicial Pra. Avg.
@@
 Table 9: Comparison of the four LLMs with different enhanced methods on eleven tasks of LexGenius, which include CoT, RAG, SFT, and GRPO.
+Appendix F With Different Enhanced Methods
+To evaluate the impact of different optimization and enhancement methods on the legal intelligence capabilities of LLMs, we selected four LLMs (including Qwen2.5-1.5B-Instruct, Qwen2.5-7B-Instruct, Qwen3-4B, and Qwen3-8B) and experimented with Supervised Fine-Tuning (SFT), Chain-of-Thought (CoT), Retrieval-Augmented Generation (RAG), and Reinforcement Learning (RL) algorithms.
+We randomly sampled 64 test instances from each of the 20 ability test sets in LexGenius, resulting in 1,280 total samples for evaluation.
+The remaining 7,105 data samples were used as the training set for SFT and RL, as well as for constructing the retrieval corpus.
+The appropriate parameters were selected for SFT and RL training.
+The experimental results of these LLMs, after applying these enhancement methods across various dimensions and tasks of legal intelligence, are shown in Table 8 and Table 9.
+Figure 30: The correlation analysis of legal intelligence ability, task, and dimension in LexGenius for 12 LLMs.
+Appendix G Correlation Analysis
+The performance of 12 LLMs on LexGenius is utilized to analyze correlations (see Figure 30).
+It illustrates that most of the legal intelligence abilities (left), tasks (upper right), and dimensions (lower right) exhibit low correlations.
+It shows the effectiveness of LexGenius because the low intercorrelation suggests LLMs cannot rely on general legal heuristics or shallow transfer across domains to perform well; instead, success in one category does not guarantee success in others.
+This reflects the comprehensive coverage and conceptual independence of our benchmark dimensions, further validating their robustness as an evaluation framework.
+Appendix H Limitations
+Although LexGenius structures legal general intelligence evaluation, it is limited by a lack of multimodal capabilities, cross-jurisdictional coverage, and temporal awareness.
+These gaps constrain its ability to capture real-world complexity.
+The following subsections detail these core limitations.
+Lack of Multimodal Tasks Limits Realistic Evidence Modeling.
+The current version of LexGenius relies entirely on pure textual materials, excluding multimodal evidence types common in real-world cases, such as scanned contracts, video stills, or audio transcriptions.
+This unimodal design fails to assess capabilities in visual perception, auditory understanding, and cross-modal reasoning essential for handling actual judicial cases.
+Consequently, the absence of multimodal inputs limits LLM applicability in tasks such as evidence review, fact reconstruction, and visual-legal interpretation, reducing evaluation fidelity to real-world scenarios.
+Linguistic and Jurisdictional Limitations Undermine Cross-Cultural Generalization.
+LexGenius is currently constructed solely from Chinese corpora and Mainland China’s law system, exhibiting distinct linguistic and legal singularity.
+Consequently, evaluations are confined to this context, failing to capture broader capabilities like interpreting international statutes or comparative analysis.
+This restriction limits applicability in global legal services and cross-border disputes, hampering transferability and impeding the evolution into a universal legal intelligence system.
+Lack of Evaluation on Temporal Sensitivity and Legal Validity Awareness.
+A core characteristic of law is its temporal nature.
+Applicable rules for a given issue may vary across time, especially before and after legislative amendments.
+LexGenius currently does not incorporate a systematic temporal dimension to assess whether models can understand the time-bound applicability of statutes, the validity period of precedents, or transitional legal provisions.
+Without such temporal sensitivity tests, models may produce outdated or legally invalid answers when facing evolving legal frameworks, with no mechanism to detect these errors.
+Appendix I Future Work
+While LexGenius establishes a structured evaluation framework for Chinese legal general intelligence, it has yet to fully capture real-world complexity. Therefore, our future work focuses on:
+Incorporating Multimodal Tasks to Enhance Realistic Evidence Modeling.
+The current version of LexGenius relies solely on text and does not include multimodal information common in real legal cases, such as scanned contracts, courtroom audio, or surveillance stills.
+The absence of such inputs limits the evaluation of model capabilities in visual perception, auditory comprehension, and cross-modal reasoning, essential for evidence review, fact reconstruction, and interpretation of visual-legal content.
+In the future, we plan to embed images, audio, and other modalities into tasks to assess reasoning capabilities based on heterogeneous, multi-source information, thus aligning evaluation more closely with practical judicial needs.
+Expanding Linguistic and Jurisdictional Coverage to Improve Cross-Cultural Generalization.
+The current dataset is grounded in Chinese texts and the law system of Mainland China, exhibiting limitations in language and legal tradition.
+This restricts evaluation applicability.
+Future versions will incorporate texts from Hong Kong, Macau, and Taiwan, as well as English statutes and case law from common law systems.
+We aim to construct bilingual QA pairs, translation tasks, and comparative analyses to evaluate models’ capabilities in understanding, aligning, and adapting across legal and linguistic contexts.
+This expansion contributes to benchmarking models for global legal services.
+Introducing Dynamic Testing of Legal Temporality and Time Sensitivity.
+Legal applicability is highly time-dependent.
+Legal amendments can lead to different rulings, and precedents often carry specific periods of validity and applicability.
+Currently, LexGenius lacks a systematic temporal dimension, making it difficult to evaluate whether a model can identify the applicable time windows of statutes, conditions for transitional provisions, or conflicts between old and new laws.
+Future versions will include temporally structured legal tasks that require models to make dynamic judgments under varying timeframes, enhancing their understanding and adaptability to evolving legal systems.